### v3

# Training and Evaluating with Human Label Variation: An Empirical Study

Abstract
Human label variation (HLV) challenges the standard assumption that a labelled instance has a single ground truth, instead embracing the natural variation in human annotation to train and evaluate models. While various training methods and metrics for HLV have been proposed, it is still unclear which methods and metrics perform best in what settings. We propose new evaluation metrics for HLV leveraging fuzzy set theory. Since these new proposed metrics are differentiable, we then in turn experiment with employing these metrics as training objectives. We conduct an extensive study over 6 HLV datasets testing 14 training methods and 6 evaluation metrics. We find that training on either disaggregated annotations or soft labels performs best across metrics, outperforming training using the proposed training objectives with differentiable metrics. We also show that our proposed soft metric is more interpretable and correlates best with human preference.111Code is available at: https://github.com/kmkurn/train-eval-hlv
1 Introduction
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth (Plank, 2022). Without this assumption, model training and evaluation are no longer straightforward. For example, standard metrics such as accuracy assume a single ground truth label exists for each instance. To address this issue, recent work (Pavlick and Kwiatkowski, 2019; Peterson et al., 2019; Nie, Zhou, and Bansal, 2020; Uma et al., 2020, 2021; Cui, 2023; Gajewska, 2023; Maity et al., 2023; Wan and Badillo-Urquiola, 2023, inter alia) has proposed various methods to train models and metrics to evaluate their performance in the HLV context. However, the lack of a systematic study means it is still unclear which methods and metrics perform best in what settings.
In natural language processing (NLP), most existing metrics for HLV represent human judgements as probability distributions over classes and employ information-theoretic measures such as divergence. It has largely ignored the field of remote sensing that has dealt with issues in model evaluation against fuzzy ground truths (Foody, 1996; Binaghi et al., 1999; Lewis and Brown, 2001; Pontius and Cheuk, 2006; Pontius and Connors, 2006; Silván-Cárdenas and Wang, 2008; Gómez, Biging, and Montero, 2008). In remote sensing, the task is to classify patches of a satellite image of a region into categories such as land, water, and so forth. As an image patch often contains multiple categories, the field has developed evaluation metrics that can handle this variation in the ground truth labels. Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. To the best of our knowledge, we are the first to propose employing soft metrics for HLV in NLP.
Because these soft metrics are differentiable, we next explore the use of these soft metrics as training objectives. Through an extensive study testing 14 training methods and 6 evaluation metrics on 6 HLV datasets, we show that training on either disaggregated annotations or soft labels often achieves the best performance across metrics. We also conduct a meta-evaluation on the evaluation metrics. Albeit modest, we find that our proposed soft metric correlates best with human preference. However, this suggests that there is still substantial future work to develop better evaluation metrics for HLV.
To summarise, we: (a) propose soft metrics to measure model performance in the context of HLV, taking inspiration from remote sensing research; (b) propose new training methods based on the soft evaluation metrics; and (c) conduct an extensive study covering 6 HLV datasets, 14 training methods and 6 evaluation metrics to understand the best training methods and evaluation metrics.
2 Related Work
Training methods for human label variation
First introduced by Plank (2022), the term human label variation (HLV) captures the fact that disagreeements in annotations can be well-founded and thus signal for data-driven methods. It challenges the traditional notion in machine learning that an instance has a single ground truth. Training methods that accommodate such variation had been proposed before the term was coined (Sheng, Provost, and Ipeirotis, 2008; Peterson et al., 2019; Uma et al., 2020, inter alia). Newer methods have also been proposed in the literature (Deng et al., 2023; Lee, An, and Thorne, 2023; Chen et al., 2024; Rodríguez-Barroso et al., 2024) and in the Learning with Disagreement shared task (Leonardelli et al., 2023) which provides a benchmark for HLV. While a systematic investigation of some of these methods exists (Uma et al., 2020), it employs smaller pretrained language models and covers only binary or multiclass classification tasks. In contrast, our work additionally employs a large language model and covers multilabel tasks. Furthermore, we also propose new soft evaluation metrics for HLV.
Evaluation of soft classification in remote sensing
While evaluating in the HLV context is a relatively new concept in NLP, evaluating against a fuzzy222The terms “fuzzy” and “crisp” are sometimes used in remote sensing literature to mean “soft” and “hard” in NLP literature. reference is a well-studied area in remote sensing research. Early work by Foody (1996) used a measure of distance that is equivalent to twice the Jenson-Shannon divergence (Lin, 1991). Binaghi et al. (1999) argued that entropy-based measures are sensible only if the reference is crisp, and proposed the use of fuzzy set theory to compute the soft version of standard evaluation metrics such as accuracy. A similar approach was also proposed by Harju and Mesaros (2023) for audio data. Key to this approach is the minimum function used to compute the class membership scores given a fuzzy reference and a fuzzy model output. Other proposed functions include product (Lewis and Brown, 2001), sum (Pontius and Connors, 2006), and a composite of such functions designed to ensure diagonality of the confusion matrix when the reference and the model output match perfectly (Pontius and Cheuk, 2006; Silván-Cárdenas and Wang, 2008). Approaches that take the cost of misclassifications into account have also been proposed (Gómez, Biging, and Montero, 2008). In this work, we take the fuzzy set approach by Binaghi et al. (1999) and apply it to text classification tasks.
3 Evaluation Metrics for HLV
We now discuss the evaluation metrics relevant for HLV data, first on multiclass metrics (Section 3.1), then multilabel metrics (Section 3.2), and finally the relationship between these metrics (Section 3.3). In these discussions, we will also introduce our use of soft metrics and explain how they relate to their hard variant counterparts. Table 1 provides a summary of the existing and proposed/new evaluation metrics.
Notation
Let denote the true human judgement for example and class , i.e. proportion of humans labelling example with class . Let denote the value of predicted by a model. An evaluation metric is a function such that the scalar measures how well (i.e., a positive orientation) captures the judgement distribution in . To complete our notation, let and denote the number of examples and classes respectively.
| Multiclass | Multilabel | |
| Existing: | Accuracy | Micro F1 |
| Macro F1 | Macro F1 | |
| PO-JSD | ||
| Entropy correlation | ||
| Proposed: | Soft accuracy | Soft micro F1 |
| Soft macro F1 | Soft macro F1 | |
| Multilabel PO-JSD | ||
| Multilabel entropy correlation |
3.1 Multiclass Metrics
In multiclass classification tasks, an example can only be assigned to one class. This is reflected by the unity constraints and for all . We study a total of 6 multiclass evaluation metrics . The first two are standard hard metrics, the next two are their soft versions that we propose inspired by prior work (Binaghi et al., 1999; Harju and Mesaros, 2023), and the last two are existing HLV metrics.
Let and denote and respectively. The 6 metric definitions under our notation are as follows:
1.
(Hard) accuracy, where:
This is the standard evaluation metric for classification tasks.
2.
(Hard) macro F1, where
This metric addresses the issue with accuracy that is biased toward the majority class when class proportions are imbalanced.
3.
Soft accuracy, where:
This metric is novel to this work, for evaluation in the context of HLV. This is a special case of the soft micro F1 score proposed later in Section 3.2.
4.
Soft macro F1, where:
This metric is also novel to this work, to address the issue of class imbalance with soft accuracy and other existing HLV metrics.333In the context of HLV, we understand the concept of class imbalance analogously to that in the context where class assignments are hard: there exists a class such that for all . Note that this metric implies the existence of soft class-wise metrics (see Appendix).
5.
Jensen-Shannon divergence (Uma et al., 2021), which we modify into its positively oriented version (PO-JSD for short) where:
Vectors denote the -th row of respectively. The scalar is the Jensen-Shannon divergence (Lin, 1991), where is the Kullback-Leibler divergence.444Jensen-Shannon divergence of two distributions has an upper bound of if the logarithms used in are of base . Normalising this bound to 1 results in logarithms of base 2 as .
6.
Entropy correlation (Uma et al., 2020), where:
Scalar is the normalised entropy of row of , and is the mean of . Scalars and are defined analogously.
By noting that , soft accuracy can be interpreted as the proportion of the judgement distribution that is correctly predicted. This interpretation is similar to the hard counterpart (i.e. proportion of examples that are correctly predicted) and arguably more intuitive than that of existing metrics because it is stated directly in terms of the problem of predicting human judgement distribution. Other similarities include symmetry (due to the symmetry of ) and boundedness (between 0 and 1), where the lower and the upper bounds are achieved if and only if and respectively for all . Note also that soft accuracy is reduced to the standard hard accuracy when all rows of both and are one-hot vectors (i.e. no label variation).
While functions other than have been used to compute soft accuracy (see Section 2), they lead to the soft accuracy being in the case where . This is problematic because intuitively, good HLV evaluation metrics should produce the highest score when the predicted judgement distribution perfectly matches the true counterpart. Therefore, we use the function in this work.
Other evaluation metrics we considered included entropy similarity (Uma et al., 2021), (negative) cross-entropy (Peterson et al., 2019; Pavlick and Kwiatkowski, 2019), (negative) Kullback-Leibler divergence (Nie, Zhou, and Bansal, 2020), (the positively oriented version of) information closeness (Foody, 1996) and (the positively oriented version of) Jensen-Shannon distance (Nie, Zhou, and Bansal, 2020). However, we ultimately excluded them from this study because of their expected high correlation with either PO-JSD or entropy correlation.
3.2 Multilabel Metrics
Next, we consider the case where the unity constraints do not hold, i.e. an example can be (soft) assigned to multiple classes, forming a multilabel classification task.
By considering this case as multiple independent binary classification tasks, some existing metrics can be extended straightforwardly by taking the average of the metric values over classes. For example, PO-JSD can be modified into the multilabel version:
where is a judgement distribution over two classes, and is defined analogously. Entropy correlation can also be modified similarly into the multilabel version:
where , , , and scalars , , and are defined analogously.
While it is possible to define (hard) accuracy for this multilabel case, a more common evaluation metric is the micro F1 score, which in our notation can be expressed as:
where , and is defined analogously. Similarly, we define the soft version of the micro F1 score as follows (see derivations in the Appendix):
The (hard) micro F1 score is a special case of this soft counterpart when and (i.e. no label variation). Furthermore, the soft accuracy in Section 3.1 is a special case of this soft micro F1 score when the unity constraints hold. Therefore, the soft micro F1 score is a general evaluation metric for classification tasks.
3.3 Relationship between Metrics
To analyse the relationship between metrics, we compute Pearson correlation between each pair of multiclass metrics (excluding the macro ones) similar to prior work (Chicco and Jurman, 2020). For a given number classes, we sample matrices and and compute the value of each metric on these pairs of and . The Pearson correlations between pairs of metrics are then computed based on these data points. We set for computational reasons. We report the correlation coefficients in Table 2.
Table 2 shows that the metrics are generally weakly correlated with each other, with the exception of soft accuracy and PO-JSD where the two are consistently strongly correlated (). This strong correlation with an established HLV metric gives an empirical assurance that our proposed soft accuracy is a sensible metric. The table also shows that soft accuracy is only moderately correlated with (hard) accuracy in most cases, where this correlation gets weaker when the number of classes is large (), or either or is dense ( or ). This finding suggests that hard and soft accuracy capture different aspects when humans disagree. Furthermore, the table shows that entropy correlation is poorly correlated with all metrics, which suggests that it is an outlier HLV metric.
Soft Accuracy and PO-JSD
| A-J | A-E | A-S | J-E | J-S | E-S | |
| 10 | ||||||
| 100 | ||||||
| 10 | ||||||
| 100 | ||||||
| 10 | ||||||
| 100 | ||||||
| 10 | ||||||
| 100 | ||||||
Despite their strong positive correlation, soft accuracy and PO-JSD have a notable difference: soft accuracy is upper-bounded by PO-JSD.
Theorem 3.1.
Consider the same definitions of and used in Section 3. The soft accuracy and the PO-JSD between and satisfy the following inequality:
where denote the Jensen-Shannon divergence between the -th rows of and .
Proof 3.2.
Given a scalar and discrete distributions , it has been shown that the following bound holds (Lin, 1991, Theorem 4):
where is the entropy of , and is the general form of Jensen-Shannon divergence with two distributions. It can be easily shown that is reduced to the standard Jensen-Shannon divergence when . Performing this substitution and noting that ,555All logarithms are of base 2. we have
for all . Doubling both sides and taking the average of this inequality over give the desired result.
This fact implies that soft accuracy penalises a model more heavily than PO-JSD, as illustrated in Figure 1. It shows that a skewed predicted judgement distribution of has a soft accuracy of , much less than the PO-JSD (). The high PO-JSD is potentially misleading as it is very close to the maximum possible value of 1.
4 Training Methods for HLV
| Name | A | M | L |
| Repeated labelling (ReL) | ✗ | ✓ | ✓ |
| Majority voting (MV) | ✗ | ✓ | ✗ |
| Annotator ranking (AR) | ✓ | ✓ | ✗ |
| Annotator ranking (hard) (ARh) | ✓ | ✓ | ✗ |
| Ambiguous labelling (AL) | ✗ | ✗ | ✓ |
| Soft labelling (SL) | ✗ | ✓ | ✗ |
| Multitask of SL and MV (SLMV) | ✗ | ✓ | ✗ |
| Annotator ensemble (AE) | ✓ | ✓ | ✗ |
| Annotator ensemble (hard) (AEh) | ✓ | ✓ | ✗ |
| Below are proposed in this work | |||
| Jensen-Shannon divergence (JSD) | ✗ | ✓ | ✗ |
| Soft micro F1 (SmF1) | ✗ | ✓ | ✗ |
| Soft macro F1 (SMF1) | ✗ | ✓ | ✗ |
| Loss aggregation with (LA-min) | ✗ | ✓ | ✗ |
| Loss aggregation with (LA-max) | ✗ | ✓ | ✗ |
We now discuss the training methods. In total, we experiment with 14 HLV training methods, as summarised in Table 3. The first method (ReL) is the simplest and uses standard cross-entropy loss as the training objective based on prior work (Sheng, Provost, and Ipeirotis, 2008). The next 8 methods (MV, AR, ARh, AL, SL, SLMV, AE, AEh) are from the Learning with Disagreements shared task (Leonardelli et al., 2023) which also use standard cross-entropy loss. The next 3 methods (JSD, SmF1, SMF1) use soft metrics that are differentiable as the training objective. The last 2 methods (LA-min, LA-max) aggregate the cross-entropy losses of all annotations for a given input. In detail:
1.
ReL (repeated labelling) trains on the disaggregated labels directly (Sheng, Provost, and Ipeirotis, 2008). For example, if the training data only has a single instance with annotations and then ReL constructs new training data .
2.
MV trains on the majority-voted labels. This is the baseline model given by the shared task organisers.
3.
AR (annotator ranking) and ARh (annotator ranking hard) weight each training instance equal to the sum of ranking scores of its annotators (Cui, 2023). Suppose that training instance is labelled as A, A, B, and A by 4 annotators; and is labelled as A, B, and B by only the first 3 annotators. Then, to compute the ranking score of an annotator:
AR computes the average judgement of the majority-voted label given by the annotator. The annotator ranking scores are (in order) , , , and .
ARh computes the average number of times the annotator agrees with the majority. The scores are , , , and .
For both methods, the ranking score is zero if the annotator never agrees with the majority. Given an instance, the training objective is maximising the probability of its majority-voted label(s), scaled by its weight as defined above.
4.
AL (ambiguous labelling) labels ambiguous instances, i.e. instances where annotators are not unanimous in their judgement, with all selected classes exactly once (Gajewska, 2023). Each instance-label pair is included in the training set. For example, if training instance is labelled as , , and by 3 annotators, and is labelled as , , and , then AL produces as the training data. Because it was developed for single-label tasks, we exclude this method from multilabel tasks.666Extending this method to multilabel tasks is beyond the scope of this work.
5.
SL (soft labelling) trains on soft labels as targets (Maity et al., 2023; Wan and Badillo-Urquiola, 2023). For example, if an instance is labelled as positive by 80% of annotators then the log-likelihood is , where and denote the predicted probability of the positive and the negative class respectively.
6.
SLMV performs multi-task learning using both the soft (SL) and the majority-voted (MV) labels as targets (Grötzinger, Heuschkel, and Drews, 2023). Two models with a shared encoder are trained to predict the two types of labels respectively, which is similar to the method of Fornaciari et al. (2021). For example, if an instance is labelled as positive by 80% of annotators then the multitask log-likelihood is , where and denote the parameters of the first and the second model respectively. We predict only the soft labels for evaluation.
7.
AE (annotator ensembling) and AEh (annotator ensembling hard) model each annotator separately followed by ensembling (Sullivan, Yasin, and Jacobs, 2023; Vitsakis et al., 2023). In this approach, we have as many models as there are annotators. Each model is trained on the labels given by a distinct annotator. In practice, we share the encoder parameters across models. Suppose that there are 3 annotators, and the predicted judgements of the positive class for an instance are 60%, 30%, and 90%. To ensemble these predictions at test time:
AE computes a simple mean (Sullivan, Yasin, and Jacobs, 2023). The final predicted judgement for the positive class is .
AEh computes the distribution of most likely labels (Vitsakis et al., 2023). The final predicted judgement for the positive class is because 2 out of 3 judgements are over 50%.
8.
JSD, SmF1 (soft micro F1), and SMF1 (soft macro F1) train using JSD, soft accuracy for multiclass or soft micro F1 for multilabel tasks, and soft macro F1 as the objective respectively (see Section 3 for definitions).777We exclude entropy correlation because it doesn’t have a unique maximiser. These approaches leverage the differentiability of soft metrics, resulting in a single objective for both training and inference. By directly optimising the evaluation metric at training time, we expect the models to exhibit superior performance, especially when evaluated with the corresponding metric.
9.
LA-min (loss aggregation ) and LA-max (loss aggregation ) aggregate the losses of all annotations for a given training instance using the and the functions respectively. For example, given a training instance with annotations and , we compute the losses and where is the cross-entropy loss function. Then, we update model parameters based on the gradient of where is either the or the function. Using the (resp. ) function means selecting the least (resp. most) “surprising” annotation for the model. Using the mean aggregation is mathematically equivalent to SL and thus excluded.
5 Evaluation of HLV Training Methods
We now conduct an extensive study covering 6 datasets, 14 training methods, and 6 evaluation metrics across 2 language models to understand the performance landscape of HLV data.
5.1 Experimental Setup
| Name | Labels | L | O | A | M | |||
| HS-Brexit (Akhtar, Basile, and Patti, 2021) | Hate speech | EN | E | 1 | 2 | ✓ | ✗ | |
| MD-Agreement (Leonardelli et al., 2021) | Offensiveness | EN | C | 10 | 2 | ✓ | ✗ | |
| ArMIS (Almanea and Poesio, 2022) | Misogyny | AR | E | 1 | 2 | ✓ | ✗ | |
| ChaosNLI (Nie, Zhou, and Bansal, 2020) | NLI | EN | C | 3 | 3 | ✗ | ✗ | |
| MFRC (Trager et al., 2022) | Moral sentiment | EN | E | 18 | 8 | ✓ | ✓ |
Public datasets
We use 5 public datasets with disaggregated annotations that have been used in previous HLV studies, four of which are in English and one in Arabic. Table 4 shows that the datasets cover both crowd (domain non-specialist) and expert (domain specialist) annotators, as well as the number of examples, average number of annotations, and the language of the dataset. For HS-Brexit, MD-Agreement, and ArMIS, we use the same train–test splits as Leonardelli et al. (2023). For both ChaosNLI and MFRC, we use 10-fold cross-validation with 10% of the training portion used as a development set. For ChaosNLI, we include only the SNLI and MNLI subsets (each has 1.5K examples) because of their standard NLI setup of one premise and one hypothesis.
“Text Annotation Game” (TAG) dataset
We also include a dataset called TAG which consists of English texts describing legal problems written by non-expert legal help seekers. We have access to this dataset because of our collaboration with Justice Connect,888https://justiceconnect.org.au an Australian public benevolent institution999As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile providing free legal assistance to laypeople facing legal problems. Each text in the dataset was annotated by an average of 5.5 practising lawyers, who each selected one or more out of the 33 areas of law101010Including a special label “Not a legal issue”. that applied to the problem. Thus, it is a multilabel classification task. Example areas of law include Neighbourhood disputes and Housing and residential tenancies. Average inter-annotator agreement () over the areas of law is , which is modest.111111We compute this agreement on random 10% selection of the data due to the high computational cost. This is because lawyers often have different interpretations of the same problem due to their different legal specialisations and years of experience. However, these different interpretations are all valid as human label variation because these registered lawyers are subject matter experts who play a crucial part in interpreting the law. The dataset has a total of 11K texts collected between July 2020 and early December 2023. We randomly split the dataset 8:1:1 for the training, development, and test sets respectively. While it cannot be released publicly because it is based on real-world confidential legal requests from help-seekers, we include the TAG dataset due to its importance for the meta-evaluation later in Section 6.
Models
We experiment with both small and large language models. Because HS-Brexit, MD-Agreement, and ArMIS use Twitter as source data, we use TwHIN-BERT (Zhang et al., 2023) which was pretrained on Twitter data and supports both English and Arabic. For other datasets, we use RoBERTa (Liu et al., 2019). In addition, we experiment with 8B LLaMA 3 (Grattafiori et al., 2024) for the English datasets.121212While LLaMA 3 was trained on multilingual data, its intended use is English only. We replace the output layer of these pretrained models with a linear layer with output units. We use the base versions of both RoBERTa and TwHIN-BERT (with roughly 100M parameters each).
Training
We implement all methods using FlairNLP (Akbik et al., 2019). For RoBERTa and TwHIN-BERT models, we tune the learning rate and the batch size using random search. Because there are multiple evaluation metrics that aren’t necessarily comparable, we select the best hyperparameters based on the geometric mean of their values131313We transform entropy correlation to a value between 0 and 1 before taking the mean. to avoid biasing towards one metric. We exclude both hard and soft macro F1 scores from this mean as most datasets are balanced. For LLaMA, we use FlairNLP’s default hyperparameters141414Learning rate and batch size are and 32 respectively. for computational reasons. We use LoRA (Hu et al., 2022) to finetune LLaMA efficiently. All models are finetuned for 10 epochs. We truncate inputs longer than 512 tokens in the TAG dataset, affecting only 4% of instances. For ChaosNLI and MFRC, we train the final model used for evaluation of each fold on the concatenation of the training and the development sets.
Evaluation
When cross-validation is used, we evaluate the concatenated test predictions. Otherwise, we evaluate on the test set. We perform the evaluation across 3 training runs.
5.2 Results
We compute the difference between the performance of a single run of an HLV training method and the mean performance of MV. We then report the mean differences across runs. We present the results on ArMIS in Figure 2 (the only non-English dataset), HS-Brexit and MD-Agreement in Figure 3 (Learning with Disagreements shared task datasets), and the remaining datasets in Figure 4 (multiclass and multilabel datasets). We report performance of MV in Table 7 in the Appendix. We make the following observations.
First, HLV training methods have different effectiveness compared to MV, with some methods performing very poorly. For example, on ArMIS and both portions of ChaosNLI, LA-min consistently performs worse than MV. This finding suggests the importance of choosing the right training method when embracing HLV.
Second, ReL and SL substantially outperform MV on most datasets across models and evaluation metrics. Furthermore, ReL and SL are often among the top-performing methods, even when compared to JSD and SmF1 evaluated using PO-JSD and soft accuracy/micro F1 respectively. In other words, ReL and SL are often on par with (or outperform in some cases with LLaMA) the training method whose learning objective is the same as the evaluation metric, demonstrating the strength of ReL and SL.
Third, ReL and SL are generally competitive with each other. However, there are two exceptions to this trend. First, on ArMIS, SL outperforms ReL. Second, on both HS-Brexit and the SNLI portion of ChaosNLI, ReL outperforms SL with LLaMA. For the former finding, however, ArMIS is an outlier dataset: it is the only Arabic dataset, has the smallest number of examples and annotations per example, and all methods generally have higher variance on this dataset. The latter finding suggests that ReL is better suited to large language models than SL.
Fourth, SMF1 and JSD generally also perform well, although they lag behind ReL and SL. Because both SMF1 and JSD use a soft evaluation metric as the training objective, this finding underlines the value in using differentiable metrics for HLV model training.
We also observe that SmF1 with TwHIN-BERT results in undefined entropy correlation on HS-Brexit. Looking closely, we find that the method degenerates to predicting for all . We observe the same degenerate results for AEh with RoBERTa on the TAG dataset. Moreover, SmF1 with RoBERTa attains much lower soft micro F1 score than the top-performing ReL on TAG, even though SmF1 optimises the evaluation metric directly during training. Our investigation suggests that these unexpected observations are due to the class imbalance present in both HS-Brexit and TAG. When the training objective is fair (e.g., SMF1), the problem disappears.
5.3 Discussion
ReL and SL appear to perform best as a training method, and that is somewhat surprising because they are the most straightforward methods to incorporate HLV in model training. That said, our finding that SL is a strong performer is in line with what Uma et al. (2021) discovered. The results of ReL, however, stand in contrast with prior work (Uma et al., 2021; Kurniawan et al., 2024). This difference may be due to the structured prediction tasks that they considered which have an exponentially large output space. We note that Uma et al. (2021) also found that ReL performs really well in simple, multiclass classification tasks, in line with our findings.
Moreover, the results suggest that ReL is likely to outperform SL when large language models are used. A possible explanation is that with SL, the model observes the complete judgement distribution of an instance as target. Thus, there is not much flexibility as the model has to predict that distribution to minimise the training loss. In contrast, with ReL, the model only observes one annotation of the instance as target. Therefore, it has more flexibility in how it distributes the judgement distribution mass across all the labels of that instance over training steps. We hypothesise that large language models have sufficient learning capacity to benefit from this flexibility while smaller models do not. This explanation is consistent with our findings on ChaosNLI that ReL has much larger gains compared to SL. ChaosNLI has 100 annotations per instance, the largest number out of all datasets. We leave further investigation of this hypothesis to future work.
Despite its strengths, ReL has substantial computational cost: because it keeps the annotations disaggregated, the size of the training data can be enormous. For example, in the ChaosNLI dataset, ReL makes the training data 100 times larger because there are exactly 100 annotations for each training instance. That ReL performs better with LLaMA than RoBERTa on ChaosNLI exacerbates the computational inefficiency of ReL: both the data and the model must be large for it to perform optimally. Mitigating this inefficiency can be an interesting avenue for future work.
Because the training data is much larger, one may think that the superiority of ReL over SL is due to its much longer training. However, this is not the case. We find that increasing the number of training iterations of SL to match that of ReL does not improve performance.
6 Meta-Evaluation of HLV Evaluation Metrics
In the previous section, we gained an understanding about the best training methods that perform consistently well across metrics. In this section, we tackle the question of what the best evaluation metric is for HLV data. We address this question by conducting a meta-evaluation of all the evaluation metrics in question. Intuitively, a good evaluation metric should correlate well with human preference. That is, if a model has a high performance according to a certain metric, we expect human judges to prefer that model’s predictions. This forms the basis of our meta-evaluation experiment.
To obtain a reliable estimate of human preference, we rely on subject matter experts. This is so that when they disagree we can be confident that the disagreement is a result of genuine variation rather than noise. In our case, we have access to practising lawyers who are more than qualified for our task (i.e. legal area classification). As such, we use the TAG dataset to conduct our meta-evaluation experiments despite the inability to distribute it.
6.1 Experimental Setup
For each pair of pretrained model and training method, we randomly sample an average of 28 test instances from TAG. For each instance, we (a) show the instance text, the ground truth label distribution and the predicted label distribution from one of our trained models, and then (b) ask lawyers which distribution151515Because it is multilabel task, the distribution doesn’t sum to one. is more accurate. The lawyers can choose exactly one, both, or neither of the visualisations of the distributions. On average, each instance is annotated by 3.8 lawyers. To make sure the lawyers form their own expectation on the areas of law relevant to an instance, we first show the instance text and ask them to annotate it (Figure 5). After completing this task, we then show them the label distributions from the model and ground truth (Figure 6). We conduct several pilot studies with the lawyers to refine the interface design to ensure the task is clear to our target annotators.
It is worth noting that the meta-evaluation task inherently has HLV as there is not a single, objective ground truth. Every lawyer has their own preference, and our goal is to capture the collective preference of the lawyers (i.e. the proportion of lawyers who prefer the predicted distributions). To minimise the risk of lawyers figuring out which of the two judgement distribution options is the ground truth distribution, we randomise the options and the order of the instances given to the lawyers. To filter out low quality predictions, we only show areas of law whose judgement exceeds a threshold of .161616We arrive at this threshold by choosing a value that results in similar numbers of labels whose judgement exceeds the value in the predicted and the ground truth distributions.
6.2 Results
We meta-evaluate HLV evaluation metrics by computing the correlation between their value and the fraction of times the lawyers select the predicted judgement distribution. Figure 7 shows that all metrics exhibit a linear relationship with human preference, which is positive for all metrics other than entropy correlation, suggesting that it is a poor HLV metric.
Table 5 confirms that all metrics are positively correlated with human preference except for entropy correlation. Furthermore, it shows that only soft micro F1 and PO-JSD correlate with human judgements at a level of statistical significance. This finding suggests that both metrics are good for evaluation in the HLV context.
6.3 Method ranking based on human preference
Given the human annotations above, we can compute the performance of an HLV training method based on the number of times humans select the predicted judgement distribution. Table 6 shows that the rankings of the methods have some similarities to TAG results in Figures 2, 3 and 4. For instance, ReL, JSD and SL are generally strong performers (exception: SL with LLaMA). That said, we also see some discrepancies. For example, AR and LA-min come out on top here but this is not reflected in Figures 2, 3 and 4. We contend, however, that the meta-evaluation study is ultimately based on one dataset and so we should interpret these results cautiously.
| Type | Metric | Pearson |
| Hard | Micro F1 | |
| Macro F1 | ||
| Soft | Entropy correlation | |
| PO-JSD | ∗ | |
| Soft micro F1 | ∗ | |
| Soft macro F1 |
| RoBERTa | LLaMA | ||
| Rank 1–7 | Rank 8–13 | Rank 1–7 | Rank 8–13 |
| JSD | SLMV | AR | MV |
| LA-min | SMF1 | SmF1 | SL |
| SL | ARh | LA-min | ARh |
| MV | AE | ReL | AEh |
| AR | LA-max | JSD | AE |
| ReL | AEh | SLMV | SMF1 |
| SmF1 | LA-max | ||
6.4 Discussion
Putting all the results together, PO-JSD and soft micro F1 are arguably some of the best metrics for HLV data. That said, Uma et al. (2021) argued that the value of PO-JSD is typically confined within a small range and lacks an intuitive interpretation. In contrast, soft micro F1 is more interpretable. Therefore, our general recommendation is to report both metrics, but focus on the soft micro F1 score when an accessible interpretation is important (e.g., communicating with non-technical people) or one is restricted to a single metric (e.g., in hyperparameter tuning).
7 Conclusions
We propose new evaluation metrics for evaluating model predictions with human label variation (HLV). Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. Therefore, our proposed metrics have intuitive interpretations and reduce to standard hard metrics if there is no label variation. We also generalise both our proposed metrics and some of existing ones to multilabel classification settings. We mathematically prove that our proposed soft accuracy metric is upper-bounded by an existing metric based on the Jensen-Shannon divergence.
Because the soft metrics are differentiable, we propose 3 new training methods for HLV using the metrics as the training objective. Additionally, we propose 2 new training methods that aggregate losses over annotations of the same instance. We test our proposed methods on 6 datasets spanning binary, multiclass, and multilabel classification tasks, as well as crowd and expert annotators. We evaluate across 2 language models and using a total of 6 HLV evaluation metrics including both existing and proposed ones. We compare against 9 existing HLV training methods and find that training on disaggregated annotations (ReL) or soft labels (SL) performs best across evaluation metrics, and that the former is likely to perform better than the latter when both the number of annotations per instance is large and a large language model is used. Our meta-evaluation of these metrics shows that our proposed soft metric correlates best with human preference. We recommend further research to include our proposed soft metric when reporting model performance in the HLV context.
Limitations
The TAG dataset is private and cannot be released publicly, which is a limitation of our work in terms of reproducibility. This is because the data is based on real-world confidential legal requests from help-seekers. More importantly, their safety and privacy is of high concern: some cases are so unique that even if you were to anonymise the cases, re-identfication may still be possible. Nevertheless, we believe our work still offers valuable scientific knowledge on the investigation of HLV training methods and evaluation metrics.
Our empirical meta-evaluation is limited to a single multilabel classification dataset. Thus, it is unclear if the findings extend to other datasets and binary or single-label tasks. Future work could expand on this limitation by experimenting with a different or diverse selection of datasets.
Derivation of Soft F1 Scores
.1 Fuzzy Sets
Below we provide a brief overview of fuzzy sets drawn from the work of Kruse et al. (2022). A fuzzy set is defined by its membership function , where represents the degree of membership of in . The cardinality of is defined as the sum of the degrees of membership of all elements of :
If and are fuzzy sets over the same universe then their intersection is a fuzzy set whose membership function is defined as
where is a function that satisfies 3 properties: commutativity, associativity, and monotonicity.171717 where The function is called a t-norm (triangular norm). A special t-norm is the function because it is also idempotent, i.e. .
.2 Soft F1 Scores
Let and denote the predicted and the true (crisp) sets of examples in class respectively. Standard precision and recall scores for class are defined as:
Embracing HLV, we view judgement distributions as degrees of memberships of fuzzy sets, each corresponds to a class, over a universe of examples. Specifically, we consider and as the true and the predicted degrees of membership of example in class . Therefore, we can define the soft precision and recall scores using fuzzy set operations as follows (Harju and Mesaros, 2023):
and
The soft F1 score is the harmonic mean between the soft precision and recall scores as usual:
Taking the mean of Equation 3 over classes results in the soft macro F1 score. To obtain the micro variant, we simply modify the sums in both the numerator and the denominator of Equations 1 and 2 to also iterate over classes to get the soft micro precision and recall scores, and then take the harmonic mean of the two scores as normal.
Performance of MV
Table 7 reports the performance of MV across datasets and evaluation metrics.
| Dataset | Model | Micro F1 | Macro F1 | Entropy corr. | PO-JSD | Soft micro F1 | Soft macro F1 |
| HS-Brexit | TwHIN-BERT | .903 .009 | .581 .093 | .470 .119 | .941 .010 | .883 .013 | .676 .071 |
| LLaMA | .950 .012 | .660 .104 | .330 .065 | .939 .002 | .886 .003 | .623 .012 | |
| MD-Agr | TwHIN-BERT | .809 .003 | .779 .002 | .379 .002 | .921 .001 | .811 .002 | .790 .003 |
| LLaMA | .810 .001 | .784 .001 | .215 .019 | .863 .003 | .770 .002 | .743 .003 | |
| ArMIS | TwHIN-BERT | .713 .014 | .705 .014 | .062 .015 | .769 .011 | .702 .009 | .695 .008 |
| C-SNLI | RoBERTa | .632 .005 | .551 .007 | .123 .014 | .830 .005 | .672 .004 | .629 .007 |
| LLaMA | .601 .010 | .551 .008 | .118 .022 | .830 .005 | .668 .005 | .633 .007 | |
| C-MNLI | RoBERTa | .496 .014 | .369 .014 | .044 .007 | .854 .009 | .668 .010 | .642 .009 |
| LLaMA | .514 .017 | .439 .010 | .100 .018 | .827 .005 | .643 .006 | .615 .006 | |
| MFRC | RoBERTa | .657 .001 | .456 .002 | .404 .002 | .954 .000 | .591 .001 | .418 .000 |
| LLaMA | .649 .001 | .454 .005 | .353 .005 | .948 .000 | .576 .001 | .395 .001 | |
| TAG | RoBERTa | .742 .002 | .610 .014 | .394 .002 | .985 .000 | .645 .002 | .471 .003 |
| LLaMA | .759 .006 | .681 .005 | .279 .010 | .983 .000 | .648 .002 | .521 .003 |
Acknowledgements.
This research is supported by the Australian Research Council Linkage Project LP210200917181818https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution.191919As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.References
- Akbik et al. (2019) Akbik, Alan, Tanja Bergmann, Duncan Blythe, Kashif Rasul, Stefan Schweter, and Roland Vollgraf. 2019. FLAIR: An easy-to-use framework for state-of-the-art NLP. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics (Demonstrations), pages 54–59.
- Akhtar, Basile, and Patti (2021) Akhtar, Sohail, Valerio Basile, and Viviana Patti. 2021. Whose Opinions Matter? Perspective-aware Models to Identify Opinions of Hate Speech Victims in Abusive Language Detection. CoRR, cs.CL/2106.15896.
- Almanea and Poesio (2022) Almanea, Dina and Massimo Poesio. 2022. ArMIS - The Arabic Misogyny and Sexism Corpus with Annotator Subjective Disagreements. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 2282–2291.
- Binaghi et al. (1999) Binaghi, Elisabetta, Pietro A. Brivio, Paolo Ghezzi, and Anna Rampini. 1999. A fuzzy set-based accuracy assessment of soft classification. Pattern Recognition Letters, 20(9):935–948.
- Chen et al. (2024) Chen, Beiduo, Xinpeng Wang, Siyao Peng, Robert Litschko, Anna Korhonen, and Barbara Plank. 2024. “Seeing the Big through the Small”: Can LLMs Approximate Human Judgment Distributions on NLI from a Few Explanations? In Findings of the Association for Computational Linguistics: EMNLP 2024, pages 14396–14419.
- Chicco and Jurman (2020) Chicco, Davide and Giuseppe Jurman. 2020. The advantages of the Matthews correlation coefficient (MCC) over F1 score and accuracy in binary classification evaluation. BMC Genomics, 21(1):6.
- Cui (2023) Cui, Xia. 2023. Xiacui at SemEval-2023 Task 11: Learning a Model in Mixed-Annotator Datasets Using Annotator Ranking Scores as Training Weights. In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 1076–1084.
- Deng et al. (2023) Deng, Naihao, Xinliang Zhang, Siyang Liu, Winston Wu, Lu Wang, and Rada Mihalcea. 2023. You Are What You Annotate: Towards Better Models through Annotator Representations. In Findings of the Association for Computational Linguistics: EMNLP 2023, pages 12475–12498.
- Foody (1996) Foody, G. M. 1996. Approaches for the production and evaluation of fuzzy land cover classifications from remotely-sensed data. International Journal of Remote Sensing, 17(7):1317–1340.
- Fornaciari et al. (2021) Fornaciari, Tommaso, Alexandra Uma, Silviu Paun, Barbara Plank, Dirk Hovy, and Massimo Poesio. 2021. Beyond Black & White: Leveraging Annotator Disagreement via Soft-Label Multi-Task Learning. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 2591–2597.
- Gajewska (2023) Gajewska, Ewelina. 2023. Eevvgg at SemEval-2023 Task 11: Offensive Language Classification with Rater-based Information. In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 171–176.
- Gómez, Biging, and Montero (2008) Gómez, D., G. Biging, and J. Montero. 2008. Accuracy statistics for judging soft classification. International Journal of Remote Sensing, 29(3):693–709.
- Grattafiori et al. (2024) Grattafiori, Aaron, Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Alex Vaughan, Amy Yang, Angela Fan, Anirudh Goyal, Anthony Hartshorn, Aobo Yang, Archi Mitra, Archie Sravankumar, Artem Korenev, Arthur Hinsvark, Arun Rao, Aston Zhang, Aurelien Rodriguez, Austen Gregerson, Ava Spataru, Baptiste Roziere, Bethany Biron, Binh Tang, Bobbie Chern, Charlotte Caucheteux, Chaya Nayak, Chloe Bi, Chris Marra, Chris McConnell, Christian Keller, Christophe Touret, Chunyang Wu, Corinne Wong, Cristian Canton Ferrer, Cyrus Nikolaidis, Damien Allonsius, Daniel Song, Danielle Pintz, Danny Livshits, Danny Wyatt, David Esiobu, Dhruv Choudhary, Dhruv Mahajan, Diego Garcia-Olano, Diego Perino, Dieuwke Hupkes, Egor Lakomkin, Ehab AlBadawy, Elina Lobanova, Emily Dinan, Eric Michael Smith, Filip Radenovic, Francisco Guzmán, Frank Zhang, Gabriel Synnaeve, Gabrielle Lee, Georgia Lewis Anderson, Govind Thattai, Graeme Nail, Gregoire Mialon, Guan Pang, Guillem Cucurell, Hailey Nguyen, Hannah Korevaar, Hu Xu, Hugo Touvron, Iliyan Zarov, Imanol Arrieta Ibarra, Isabel Kloumann, Ishan Misra, Ivan Evtimov, Jack Zhang, Jade Copet, Jaewon Lee, Jan Geffert, Jana Vranes, Jason Park, Jay Mahadeokar, Jeet Shah, Jelmer van der Linde, Jennifer Billock, Jenny Hong, Jenya Lee, Jeremy Fu, Jianfeng Chi, Jianyu Huang, Jiawen Liu, Jie Wang, Jiecao Yu, Joanna Bitton, Joe Spisak, Jongsoo Park, Joseph Rocca, Joshua Johnstun, Joshua Saxe, Junteng Jia, Kalyan Vasuden Alwala, Karthik Prasad, Kartikeya Upasani, Kate Plawiak, Ke Li, Kenneth Heafield, Kevin Stone, Khalid El-Arini, Krithika Iyer, Kshitiz Malik, Kuenley Chiu, Kunal Bhalla, Kushal Lakhotia, Lauren Rantala-Yeary, Laurens van der Maaten, Lawrence Chen, Liang Tan, Liz Jenkins, Louis Martin, Lovish Madaan, Lubo Malo, Lukas Blecher, Lukas Landzaat, Luke de Oliveira, Madeline Muzzi, Mahesh Pasupuleti, Mannat Singh, Manohar Paluri, Marcin Kardas, Maria Tsimpoukelli, Mathew Oldham, Mathieu Rita, Maya Pavlova, Melanie Kambadur, Mike Lewis, Min Si, Mitesh Kumar Singh, Mona Hassan, Naman Goyal, Narjes Torabi, Nikolay Bashlykov, Nikolay Bogoychev, Niladri Chatterji, Ning Zhang, Olivier Duchenne, Onur Çelebi, Patrick Alrassy, Pengchuan Zhang, Pengwei Li, Petar Vasic, Peter Weng, Prajjwal Bhargava, Pratik Dubal, Praveen Krishnan, Punit Singh Koura, Puxin Xu, Qing He, Qingxiao Dong, Ragavan Srinivasan, Raj Ganapathy, Ramon Calderer, Ricardo Silveira Cabral, Robert Stojnic, Roberta Raileanu, Rohan Maheswari, Rohit Girdhar, Rohit Patel, Romain Sauvestre, Ronnie Polidoro, Roshan Sumbaly, Ross Taylor, Ruan Silva, Rui Hou, Rui Wang, Saghar Hosseini, Sahana Chennabasappa, Sanjay Singh, Sean Bell, Seohyun Sonia Kim, Sergey Edunov, Shaoliang Nie, Sharan Narang, Sharath Raparthy, Sheng Shen, Shengye Wan, Shruti Bhosale, Shun Zhang, Simon Vandenhende, Soumya Batra, Spencer Whitman, Sten Sootla, Stephane Collot, Suchin Gururangan, Sydney Borodinsky, Tamar Herman, Tara Fowler, Tarek Sheasha, Thomas Georgiou, Thomas Scialom, Tobias Speckbacher, Todor Mihaylov, Tong Xiao, Ujjwal Karn, Vedanuj Goswami, Vibhor Gupta, Vignesh Ramanathan, Viktor Kerkez, Vincent Gonguet, Virginie Do, Vish Vogeti, Vítor Albiero, Vladan Petrovic, Weiwei Chu, Wenhan Xiong, Wenyin Fu, Whitney Meers, Xavier Martinet, Xiaodong Wang, Xiaofang Wang, Xiaoqing Ellen Tan, Xide Xia, Xinfeng Xie, Xuchao Jia, Xuewei Wang, Yaelle Goldschlag, Yashesh Gaur, Yasmine Babaei, Yi Wen, Yiwen Song, Yuchen Zhang, Yue Li, Yuning Mao, Zacharie Delpierre Coudert, Zheng Yan, Zhengxing Chen, Zoe Papakipos, Aaditya Singh, Aayushi Srivastava, Abha Jain, Adam Kelsey, Adam Shajnfeld, Adithya Gangidi, Adolfo Victoria, Ahuva Goldstand, Ajay Menon, Ajay Sharma, Alex Boesenberg, Alexei Baevski, Allie Feinstein, Amanda Kallet, Amit Sangani, Amos Teo, Anam Yunus, Andrei Lupu, Andres Alvarado, Andrew Caples, Andrew Gu, Andrew Ho, Andrew Poulton, Andrew Ryan, Ankit Ramchandani, Annie Dong, Annie Franco, Anuj Goyal, Aparajita Saraf, Arkabandhu Chowdhury, Ashley Gabriel, Ashwin Bharambe, Assaf Eisenman, Azadeh Yazdan, Beau James, Ben Maurer, Benjamin Leonhardi, Bernie Huang, Beth Loyd, Beto De Paola, Bhargavi Paranjape, Bing Liu, Bo Wu, Boyu Ni, Braden Hancock, Bram Wasti, Brandon Spence, Brani Stojkovic, Brian Gamido, Britt Montalvo, Carl Parker, Carly Burton, Catalina Mejia, Ce Liu, Changhan Wang, Changkyu Kim, Chao Zhou, Chester Hu, Ching-Hsiang Chu, Chris Cai, Chris Tindal, Christoph Feichtenhofer, Cynthia Gao, Damon Civin, Dana Beaty, Daniel Kreymer, Daniel Li, David Adkins, David Xu, Davide Testuggine, Delia David, Devi Parikh, Diana Liskovich, Didem Foss, Dingkang Wang, Duc Le, Dustin Holland, Edward Dowling, Eissa Jamil, Elaine Montgomery, Eleonora Presani, Emily Hahn, Emily Wood, Eric-Tuan Le, Erik Brinkman, Esteban Arcaute, Evan Dunbar, Evan Smothers, Fei Sun, Felix Kreuk, Feng Tian, Filippos Kokkinos, Firat Ozgenel, Francesco Caggioni, Frank Kanayet, Frank Seide, Gabriela Medina Florez, Gabriella Schwarz, Gada Badeer, Georgia Swee, Gil Halpern, Grant Herman, Grigory Sizov, Guangyi, Zhang, Guna Lakshminarayanan, Hakan Inan, Hamid Shojanazeri, Han Zou, Hannah Wang, Hanwen Zha, Haroun Habeeb, Harrison Rudolph, Helen Suk, Henry Aspegren, Hunter Goldman, Hongyuan Zhan, Ibrahim Damlaj, Igor Molybog, Igor Tufanov, Ilias Leontiadis, Irina-Elena Veliche, Itai Gat, Jake Weissman, James Geboski, James Kohli, Janice Lam, Japhet Asher, Jean-Baptiste Gaya, Jeff Marcus, Jeff Tang, Jennifer Chan, Jenny Zhen, Jeremy Reizenstein, Jeremy Teboul, Jessica Zhong, Jian Jin, Jingyi Yang, Joe Cummings, Jon Carvill, Jon Shepard, Jonathan McPhie, Jonathan Torres, Josh Ginsburg, Junjie Wang, Kai Wu, Kam Hou U, Karan Saxena, Kartikay Khandelwal, Katayoun Zand, Kathy Matosich, Kaushik Veeraraghavan, Kelly Michelena, Keqian Li, Kiran Jagadeesh, Kun Huang, Kunal Chawla, Kyle Huang, Lailin Chen, Lakshya Garg, Lavender A, Leandro Silva, Lee Bell, Lei Zhang, Liangpeng Guo, Licheng Yu, Liron Moshkovich, Luca Wehrstedt, Madian Khabsa, Manav Avalani, Manish Bhatt, Martynas Mankus, Matan Hasson, Matthew Lennie, Matthias Reso, Maxim Groshev, Maxim Naumov, Maya Lathi, Meghan Keneally, Miao Liu, Michael L. Seltzer, Michal Valko, Michelle Restrepo, Mihir Patel, Mik Vyatskov, Mikayel Samvelyan, Mike Clark, Mike Macey, Mike Wang, Miquel Jubert Hermoso, Mo Metanat, Mohammad Rastegari, Munish Bansal, Nandhini Santhanam, Natascha Parks, Natasha White, Navyata Bawa, Nayan Singhal, Nick Egebo, Nicolas Usunier, Nikhil Mehta, Nikolay Pavlovich Laptev, Ning Dong, Norman Cheng, Oleg Chernoguz, Olivia Hart, Omkar Salpekar, Ozlem Kalinli, Parkin Kent, Parth Parekh, Paul Saab, Pavan Balaji, Pedro Rittner, Philip Bontrager, Pierre Roux, Piotr Dollar, Polina Zvyagina, Prashant Ratanchandani, Pritish Yuvraj, Qian Liang, Rachad Alao, Rachel Rodriguez, Rafi Ayub, Raghotham Murthy, Raghu Nayani, Rahul Mitra, Rangaprabhu Parthasarathy, Raymond Li, Rebekkah Hogan, Robin Battey, Rocky Wang, Russ Howes, Ruty Rinott, Sachin Mehta, Sachin Siby, Sai Jayesh Bondu, Samyak Datta, Sara Chugh, Sara Hunt, Sargun Dhillon, Sasha Sidorov, Satadru Pan, Saurabh Mahajan, Saurabh Verma, Seiji Yamamoto, Sharadh Ramaswamy, Shaun Lindsay, Shaun Lindsay, Sheng Feng, Shenghao Lin, Shengxin Cindy Zha, Shishir Patil, Shiva Shankar, Shuqiang Zhang, Shuqiang Zhang, Sinong Wang, Sneha Agarwal, Soji Sajuyigbe, Soumith Chintala, Stephanie Max, Stephen Chen, Steve Kehoe, Steve Satterfield, Sudarshan Govindaprasad, Sumit Gupta, Summer Deng, Sungmin Cho, Sunny Virk, Suraj Subramanian, Sy Choudhury, Sydney Goldman, Tal Remez, Tamar Glaser, Tamara Best, Thilo Koehler, Thomas Robinson, Tianhe Li, Tianjun Zhang, Tim Matthews, Timothy Chou, Tzook Shaked, Varun Vontimitta, Victoria Ajayi, Victoria Montanez, Vijai Mohan, Vinay Satish Kumar, Vishal Mangla, Vlad Ionescu, Vlad Poenaru, Vlad Tiberiu Mihailescu, Vladimir Ivanov, Wei Li, Wenchen Wang, Wenwen Jiang, Wes Bouaziz, Will Constable, Xiaocheng Tang, Xiaojian Wu, Xiaolan Wang, Xilun Wu, Xinbo Gao, Yaniv Kleinman, Yanjun Chen, Ye Hu, Ye Jia, Ye Qi, Yenda Li, Yilin Zhang, Ying Zhang, Yossi Adi, Youngjin Nam, Yu, Wang, Yu Zhao, Yuchen Hao, Yundi Qian, Yunlu Li, Yuzi He, Zach Rait, Zachary DeVito, Zef Rosnbrick, Zhaoduo Wen, Zhenyu Yang, Zhiwei Zhao, and Zhiyu Ma. 2024. The Llama 3 Herd of Models. CoRR, cs.AI/2407.21783.
- Grötzinger, Heuschkel, and Drews (2023) Grötzinger, Dennis, Simon Heuschkel, and Matthias Drews. 2023. CICL_DMS at SemEval-2023 Task 11: Learning With Disagreements (Le-Wi-Di). In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 1030–1036.
- Harju and Mesaros (2023) Harju, Manu and Annamaria Mesaros. 2023. Evaluating classification systems against soft labels with fuzzy precision and recall. In Proceedings of the 8th Detection and Classification of Acoustic Scenes and Events 2023 Workshop (DCASE2023), pages 46–50.
- Hu et al. (2022) Hu, Edward J., Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2022. LoRA: Low-Rank Adaptation of Large Language Models. In The Tenth International Conference on Learning Representations.
- Kruse et al. (2022) Kruse, Rudolf, Sanaz Mostaghim, Christian Borgelt, Christian Braune, and Matthias Steinbrecher. 2022. Introduction to Fuzzy Sets and Fuzzy Logics. In Computational Intelligence: A Methodological Introduction. Springer, pages 373–405.
- Kurniawan et al. (2024) Kurniawan, Kemal, Meladel Mistica, Timothy Baldwin, and Jey Han Lau. 2024. To Aggregate or Not to Aggregate. That is the Question: A Case Study on Annotation Subjectivity in Span Prediction. In Proceedings of the 14th Workshop on Computational Approaches to Subjectivity, Sentiment, & Social Media Analysis, pages 362–368.
- Lee, An, and Thorne (2023) Lee, Noah, Na Min An, and James Thorne. 2023. Can Large Language Models Capture Dissenting Human Voices? In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 4569–4585.
- Leonardelli et al. (2023) Leonardelli, Elisa, Gavin Abercrombie, Dina Almanea, Valerio Basile, Tommaso Fornaciari, Barbara Plank, Verena Rieser, Alexandra Uma, and Massimo Poesio. 2023. SemEval-2023 Task 11: Learning with Disagreements (LeWiDi). In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 2304–2318.
- Leonardelli et al. (2021) Leonardelli, Elisa, Stefano Menini, Alessio Palmero Aprosio, Marco Guerini, and Sara Tonelli. 2021. Agreeing to Disagree: Annotating Offensive Language Datasets with Annotators’ Disagreement. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 10528–10539.
- Lewis and Brown (2001) Lewis, H. G. and M. Brown. 2001. A generalized confusion matrix for assessing area estimates from remotely sensed data. International Journal of Remote Sensing, 22(16):3223–3235.
- Lin (1991) Lin, Jianhua. 1991. Divergence measures based on the Shannon entropy. IEEE Transactions on Information Theory, 37(1):145–151.
- Liu et al. (2019) Liu, Yinhan, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. RoBERTa: A Robustly Optimized BERT Pretraining Approach. CoRR, cs.CL/1907.11692.
- Maity et al. (2023) Maity, Ankita, Pavan Kandru, Bhavyajeet Singh, Kancharla Aditya Hari, and Vasudeva Varma. 2023. IREL at SemEval-2023 Task 11: User Conditioned Modelling for Toxicity Detection in Subjective Tasks. In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 2133–2136.
- Nie, Zhou, and Bansal (2020) Nie, Yixin, Xiang Zhou, and Mohit Bansal. 2020. What can we learn from collective human opinions on natural language inference data? In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 9131–9143.
- Pavlick and Kwiatkowski (2019) Pavlick, Ellie and Tom Kwiatkowski. 2019. Inherent Disagreements in Human Textual Inferences. Transactions of the Association for Computational Linguistics, 7:677–694.
- Peterson et al. (2019) Peterson, Joshua, Ruairidh Battleday, Thomas Griffiths, and Olga Russakovsky. 2019. Human Uncertainty Makes Classification More Robust. In 2019 IEEE/CVF International Conference on Computer Vision (ICCV), pages 9616–9625.
- Plank (2022) Plank, Barbara. 2022. The “Problem” of Human Label Variation: On Ground Truth in Data, Modeling and Evaluation. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 10671–10682.
- Pontius and Cheuk (2006) Pontius, R. G. and M. L. Cheuk. 2006. A generalized cross-tabulation matrix to compare soft-classified maps at multiple resolutions. International Journal of Geographical Information Science, 20(1):1–30.
- Pontius and Connors (2006) Pontius, R. G. and John Connors. 2006. Expanding the conceptual, mathematical, and practical methods for map comparison. In Proceedings of the Meeting of Spatial Accuracy, pages 64–79.
- Rodríguez-Barroso et al. (2024) Rodríguez-Barroso, Nuria, Eugenio Martínez Cámara, Jose Camacho Collados, M. Victoria Luzón, and Francisco Herrera. 2024. Federated Learning for Exploiting Annotators’ Disagreements in Natural Language Processing. Transactions of the Association for Computational Linguistics, 12:630–648.
- Sheng, Provost, and Ipeirotis (2008) Sheng, Victor S., Foster Provost, and Panagiotis G. Ipeirotis. 2008. Get another label? improving data quality and data mining using multiple, noisy labelers. In Proceedings of the 14th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining, pages 614–622.
- Silván-Cárdenas and Wang (2008) Silván-Cárdenas, J. L. and L. Wang. 2008. Sub-pixel confusion–uncertainty matrix for assessing soft classifications. Remote Sensing of Environment, 112(3):1081–1095.
- Sullivan, Yasin, and Jacobs (2023) Sullivan, Michael, Mohammed Yasin, and Cassandra L. Jacobs. 2023. University at Buffalo at SemEval-2023 Task 11: MASDA–Modelling Annotator Sensibilities through DisAggregation. In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 978–985.
- Trager et al. (2022) Trager, Jackson, Alireza S. Ziabari, Aida Mostafazadeh Davani, Preni Golazizian, Farzan Karimi-Malekabadi, Ali Omrani, Zhihe Li, Brendan Kennedy, Nils Karl Reimer, Melissa Reyes, Kelsey Cheng, Mellow Wei, Christina Merrifield, Arta Khosravi, Evans Alvarez, and Morteza Dehghani. 2022. The Moral Foundations Reddit Corpus. CoRR, cs.CL/2208.05545.
- Uma et al. (2020) Uma, Alexandra, Tommaso Fornaciari, Dirk Hovy, Silviu Paun, Barbara Plank, and Massimo Poesio. 2020. A Case for Soft Loss Functions. Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, 8(1):173–177.
- Uma et al. (2021) Uma, Alexandra N., Tommaso Fornaciari, Dirk Hovy, Silviu Paun, Barbara Plank, and Massimo Poesio. 2021. Learning from Disagreement: A Survey. Journal of Artificial Intelligence Research, 72:1385–1470.
- Vitsakis et al. (2023) Vitsakis, Nikolas, Amit Parekh, Tanvi Dinkar, Gavin Abercrombie, Ioannis Konstas, and Verena Rieser. 2023. iLab at SemEval-2023 Task 11 Le-Wi-Di: Modelling Disagreement or Modelling Perspectives? In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 1660–1669.
- Wan and Badillo-Urquiola (2023) Wan, Ruyuan and Karla Badillo-Urquiola. 2023. Dragonfly_captain at SemEval-2023 Task 11: Unpacking Disagreement with Investigation of Annotator Demographics and Task Difficulty. In Proceedings of the 17th International Workshop on Semantic Evaluation (SemEval-2023), pages 1978–1982.
- Zhang et al. (2023) Zhang, Xinyang, Yury Malkov, Omar Florez, Serim Park, Brian McWilliams, Jiawei Han, and Ahmed El-Kishky. 2023. TwHIN-BERT: A Socially-Enriched Pre-trained Language Model for Multilingual Tweet Representations at Twitter. In Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining, pages 5597–5607.