### v5

# Training and Evaluating with Human Label Variation: An Empirical Study

Abstract
Human label variation (HLV) challenges the standard assumption that a labelled instance has a single ground truth, instead embracing the natural variation in human annotation to train and evaluate models. While various training methods and metrics for HLV have been proposed, it is still unclear which methods and metrics perform best in what settings. We propose new evaluation metrics for HLV leveraging fuzzy set theory. Since these new proposed metrics are differentiable, we then in turn experiment with employing these metrics as training objectives. We conduct an extensive study over 6 HLV datasets testing 14 training methods and 6 evaluation metrics. We find that training on either disaggregated annotations or soft labels performs best across metrics, outperforming training using the proposed training objectives with differentiable metrics. We also show that our proposed soft micro F1 score is one of the best metrics for HLV data. [^1] Code and supplementary materials are available at: https://github.com/kmkurn/train-eval-hlv
1 Introduction
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth (Plank, 2022). With HLV, different human annotations for the same instance are viewed as useful signal rather than undesirable noise. In this context, model training and evaluation are no longer straightforward. For example, standard metrics such as accuracy assume that a single ground truth label exists for each instance. To address this issue, recent work (Pavlick and Kwiatkowski, 2019; Peterson et al., 2019; Nie, Zhou, and Bansal, 2020; Uma et al., 2020, 2021; Cui, 2023; Gajewska, 2023; Maity et al., 2023; Wan and Badillo-Urquiola, 2023, inter alia) has proposed various methods to train models and metrics to evaluate their performance in the HLV context. However, the lack of a systematic study means it is still unclear which methods and metrics perform best.
In natural language processing (NLP), most existing metrics for HLV represent human judgements as probability distributions over classes and employ information-theoretic measures such as divergence. It has largely ignored the field of remote sensing that has dealt with issues in model evaluation against fuzzy ground truths (Foody, 1996; Binaghi et al., 1999; Lewis and Brown, 2001; Pontius and Cheuk, 2006; Pontius and Connors, 2006; Silván-Cárdenas and Wang, 2008; Gómez, Biging, and Montero, 2008). In remote sensing, the task is to classify patches of a satellite image of a region into categories such as land, water, and so forth. As an image patch often contains multiple categories, the field has developed evaluation metrics that can handle this variation in the ground truth labels.
Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. In contrast to information-theoretic measures, each of these soft metrics has an intuitive interpretation that corresponds to the standard counterpart. For example, soft accuracy is interpreted as the proportion of correctly predicted judgement distributions. To the best of our knowledge, we are the first to propose employing soft metrics for HLV in NLP.
As these soft metrics are differentiable, we next explore the use of each soft metric as the training objective. We perform an extensive study testing 14 training methods and 6 evaluation metrics on 6 HLV datasets across 2 pretrained models of different sizes to investigate which training methods are best for HLV data across the different metrics. This study is then followed by an empirical meta-evaluation of the evaluation metrics to understand which evaluation metrics are best for HLV data. To the best of our knowledge, we are the first to perform such an empirical meta-evaluation of HLV evaluation metrics.
To summarise, we: (a) propose soft metrics to measure model performance in the context of HLV, taking inspiration from remote sensing research; (b) propose new training methods based on the soft evaluation metrics; and (c) conduct an extensive study covering 6 HLV datasets, 14 training methods, 2 pretrained models, and 6 evaluation metrics to understand the best training methods and evaluation metrics.
We find that two of the simplest methods for HLV training perform surprisingly well, outperforming more complex methods including training with each soft metric as the objective and attaining the best performance in most cases. The first method considers each annotation of an instance as a separate instance–label pair. Training is then performed on this disaggregated data where the same instance may occur multiple times with a different label each. The second method trains the model to predict the full label distribution induced by the instance’s annotations using the standard cross-entropy loss. This is in contrast to the standard model training where the prediction target is the mode of the label distribution.
For evaluation metrics, we find that an existing metric based on Jensen-Shannon divergence (JSD) and our proposed soft micro F1 score are two of the best HLV metrics. Our analysis shows that in single-label classification, the two metrics are highly correlated but we prove that soft accuracy (because micro F1 is equal to accuracy in this case) is upper-bounded by the JSD-based metric. As a result, we show that the JSD-based metric can give a score close to its maximum value to wrong predictions, which is potentially misleading, unlike our proposed soft accuracy. We thus recommend future work in HLV to report both metrics but focus on our proposed soft metrics when interpretability is important.
2 Related Work
Training methods for human label variation
First introduced by Plank (2022), the term human label variation (HLV) captures the fact that disagreeements in annotations can be well-founded and thus signal for data-driven methods. It challenges the traditional notion in machine learning that an instance has a single ground truth. Training methods that accommodate such variation had been proposed before the term was coined (Sheng, Provost, and Ipeirotis, 2008; Peterson et al., 2019; Uma et al., 2020, inter alia). Newer methods have also been proposed in the literature (Deng et al., 2023; Lee, An, and Thorne, 2023; Chen et al., 2024; Rodríguez-Barroso et al., 2024) and in the Learning with Disagreement shared task (Leonardelli et al., 2023) which provides a benchmark for HLV. While a systematic investigation of some of these methods exists (Uma et al., 2020), it employs smaller pretrained language models and covers only binary or multiclass classification tasks. In contrast, our work additionally employs a large language model and covers multilabel tasks. Furthermore, we also propose new soft evaluation metrics for HLV.
Evaluation of soft classification in remote sensing
While evaluating in the HLV context is a relatively new concept in NLP, evaluating against a fuzzy [^2] The terms “fuzzy” and “crisp” are sometimes used in remote sensing literature to mean “soft” and “hard” in NLP literature. reference is a well-studied area in remote sensing research. Early work by Foody (1996) used a measure of distance that is equivalent to twice the Jensen-Shannon divergence (Lin, 1991). Binaghi et al. (1999) argued that entropy-based measures are sensible only if the reference is crisp, and proposed the use of fuzzy set theory to compute the soft version of standard evaluation metrics such as accuracy. A similar approach was also proposed by Harju and Mesaros (2023) for audio data. Key to this approach is the minimum function used to compute the class membership scores given a fuzzy reference and a fuzzy model output. Other proposed functions include product (Lewis and Brown, 2001), sum (Pontius and Connors, 2006), and a composite of such functions designed to ensure diagonality of the confusion matrix when the reference and the model output match perfectly (Pontius and Cheuk, 2006; Silván-Cárdenas and Wang, 2008). Approaches that take the cost of misclassifications into account have also been proposed (Gómez, Biging, and Montero, 2008). In this work, we take the fuzzy set approach by Binaghi et al. (1999) and apply it to text classification tasks.
Meta-evaluation of HLV evaluation metrics
There is little existing work on the meta-evaluation of HLV metrics. Most studies used information theory-based measures such as cross-entropy (Uma et al., 2020; Leonardelli et al., 2023, inter alia), presumably because of the standard probabilistic approach in modern NLP. Rizzi et al. (2024) proposed some theoretical properties that an ideal soft metric should satisfy and performed a theoretical meta-evaluation of existing HLV metrics for single-label classification. In contrast, we propose new soft metrics for both single- and multilabel classification and perform an empirical meta-evaluation of these metrics.
3 Evaluation Metrics for HLV
We now discuss the evaluation metrics relevant for HLV data, first on multiclass metrics (Section˜3.1), then multilabel metrics (Section˜3.2), and finally the relationship between these metrics (Section˜3.3). In these discussions, we will also introduce our use of soft metrics and explain how they relate to their hard variant counterparts. Table˜1 provides a summary of the existing and proposed/new evaluation metrics.
Notation
Let $P_{ik}\geq 0$ denote the human judgement for example $i$ and class $k$, i.e. proportion of humans labelling example $i$ with class $k$. Let $Q_{ik}\geq 0$ denote the value of $P_{ik}$ predicted by a model. An evaluation metric $m$ is a function such that the scalar $m(\mathbf{P},\mathbf{Q})$ measures how well (i.e., a positive orientation) $\mathbf{Q}$ captures the judgement distribution in $\mathbf{P}$. To complete our notation, let $N$ and $K$ denote the number of examples and classes respectively.
Table 1: HLV metrics for classification tasks.
Multiclass Multilabel
Existing: Accuracy Micro F1
Macro F1 Macro F1
PO-JSD
Entropy correlation
Proposed: Soft accuracy Soft micro F1
Soft macro F1 Soft macro F1
Multilabel PO-JSD
Multilabel entropy correlation
3.1 Multiclass Metrics
In multiclass classification tasks, an example can only be assigned to one class. This is reflected by the unity constraints $\sum_{k}P_{ik}=1$ and $\sum_{k}Q_{ik}=1$ for all $i$. We study a total of 6 multiclass evaluation metrics $m$. The first two are standard hard metrics, the next two are their soft versions that we propose inspired by prior work (Binaghi et al., 1999; Harju and Mesaros, 2023), and the last two are existing HLV metrics.
Let $I^{P}_{ik}$ and $I^{Q}_{ik}$ denote $\mathds{1}(\arg\max_{l}P_{il}=k)$ and $\mathds{1}(\arg\max_{l}Q_{il}=k)$ respectively. The 6 metric definitions under our notation are as follows:
  1. 1.
    (Hard) accuracy, where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{N}\sum_{ik}I^{P}_{ik}I^{Q}_{ik}.$
    This is the standard evaluation metric for classification tasks.
  2. 2.
    (Hard) macro F1, where
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{K}\sum_{k}\frac{2\sum_{i}I^{P}_{ik}I^{Q}_{ik}}{\sum_{i}\left(I^{P}_{ik}+I^{Q}_{ik}\right)}.$
    This metric addresses the issue with accuracy that is biased toward the majority class when class proportions are imbalanced.
  3. 3.
    Soft accuracy, where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{N}\sum_{ik}\min(P_{ik},Q_{ik}).$
    This metric is novel to this work, for evaluation in the context of HLV. This is a special case of the soft micro F1 score proposed later in Section˜3.2.
  4. 4.
    Soft macro F1, where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{K}\sum_{k}\frac{2\sum_{i}\min(P_{ik},Q_{ik})}{\sum_{i}(P_{ik}+Q_{ik})}.$
    This metric is also novel to this work, to address the issue of class imbalance with soft accuracy and other existing HLV metrics. [^3] In the context of HLV, we understand the concept of class imbalance analogously to that in the context where class assignments are hard: there exists a class $k$ such that $\sum_{i}P_{ik}\operatorname{>\!\!>}\sum_{i}P_{il}$ for all $l\neq k$. Note that this metric implies the existence of soft class-wise metrics (see Appendix).
  5. 5.
    Jensen-Shannon divergence (Uma et al., 2021), which we modify into its positively oriented version (PO-JSD for short) where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}1-\frac{1}{N}\sum_{i}\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i}).$
    Vectors $\mathbf{p}_{i},\mathbf{q}_{i}$ denote the $i$-th row of $\mathbf{P},\mathbf{Q}$ respectively. The scalar $\operatorname{JSD}(\mathbf{a},\mathbf{b})=\frac{1}{2}\left(\operatorname{KL}(\mathbf{a},\frac{1}{2}\left(\mathbf{a}+\mathbf{b}\right))+\operatorname{KL}(\mathbf{b},\frac{1}{2}\left(\mathbf{a}+\mathbf{b}\right))\right)$ is the Jensen-Shannon divergence (Lin, 1991), where $\operatorname{KL}(\mathbf{a},\mathbf{b})=\sum_{k}a_{k}\log_{2}\frac{a_{k}}{b_{k}}$ is the Kullback-Leibler divergence. [^4] Jensen-Shannon divergence of two distributions has an upper bound of $\log_{b}2$ if the logarithms used in $\operatorname{KL}(\cdot)$ are of base $b$. Normalising this bound to 1 results in logarithms of base 2 as $\log_{b}x/\log_{b}2=\log_{2}x$.
  6. 6.
    Entropy correlation (Uma et al., 2020), where:
    $m(\mathbf{P},\mathbf{Q})$ $\overset{\mathrm{def}}{=}\frac{\sum_{i}\zeta^{P}_{i}\zeta^{Q}_{i}}{\sqrt{\sum_{i}{\zeta^{P}_{i}}^{2}}\sqrt{\sum_{i}{\zeta^{Q}_{i}}^{2}}},$
    $\zeta^{P}_{i}$ $=\eta^{P}_{i}-\bar{\eta}^{P},\text{ and }$
    $\zeta^{Q}_{i}$ $=\eta^{Q}_{i}-\bar{\eta}^{Q}.$
    Scalar $\eta^{P}_{i}=-\sum_{k}\frac{P_{ik}\log P_{ik}}{\log K}$ is the normalised entropy of row $i$ of $\mathbf{P}$, and $\bar{\eta}^{P}=\frac{1}{N}\sum_{i}\eta^{P}_{i}$ is the mean of $\{\eta^{P}_{i}\}_{i=1}^{N}$. Scalars $\eta^{Q}_{i}$ and $\bar{\eta}^{Q}$ are defined analogously.
Interpretation of soft accuracy
By noting that $N=\sum_{ik}P_{ik}$, soft accuracy can be interpreted as the proportion of the judgement distribution that is correctly predicted. For example, if soft accuracy is 70%, then the model predicts 70% of judgement distribution correctly. This interpretation is similar to the hard counterpart (i.e. proportion of examples that are correctly predicted) and arguably more intuitive than that of existing metrics because it is stated directly in terms of the problem of predicting human judgement distribution.
Besides interpretation, soft accuracy’s similarities to the hard counterpart include symmetry (due to the symmetry of $\min$) and boundedness (between 0 and 1), where the lower and the upper bounds are achieved if and only if $P_{ik}Q_{ik}=0$ and $P_{ik}=Q_{ik}$ respectively for all $i,k$. Note also that soft accuracy is reduced to the standard hard accuracy when all rows of both $\mathbf{P}$ and $\mathbf{Q}$ are one-hot vectors (i.e. no label variation).
While functions other than $\min$ have been used to compute soft accuracy (see Section˜2), they lead to the soft accuracy being <1<1 in the case where $\mathbf{P}=\mathbf{Q}$. This is problematic because intuitively, good HLV evaluation metrics should produce the highest score when the predicted judgement distribution perfectly matches the true counterpart. Therefore, we use the $\min$ function in this work.
Other evaluation metrics we considered included entropy similarity (Uma et al., 2021), (negative) cross-entropy (Peterson et al., 2019; Pavlick and Kwiatkowski, 2019), (negative) Kullback-Leibler divergence (Nie, Zhou, and Bansal, 2020), (the positively oriented version of) information closeness (Foody, 1996) and (the positively oriented version of) Jensen-Shannon distance (Nie, Zhou, and Bansal, 2020). However, we ultimately excluded them from this study because of their expected high correlation with either PO-JSD or entropy correlation.
3.2 Multilabel Metrics
Next, we consider the case where the unity constraints do not hold, i.e. an example can be (soft) assigned to multiple classes, forming a multilabel classification task.
By considering this case as multiple independent binary classification tasks, some existing metrics can be extended straightforwardly by taking the average of the metric values over classes. For example, PO-JSD can be modified into the multilabel version:
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}1-\frac{1}{NK}\sum_{ik}\operatorname{JSD}(P^{\text{bin}}_{ik},Q^{\text{bin}}_{ik})$
where $P^{\text{bin}}_{ik}=[P_{ik},1-P_{ik}]$ is a judgement distribution over two classes, and $Q^{\text{bin}}_{ik}$ is defined analogously. Entropy correlation can also be modified similarly into the multilabel version:
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{K}\sum_{k}\frac{\sum_{i}\zeta^{P}_{ik}\zeta^{Q}_{ik}}{\sqrt{\sum_{i}{\zeta^{P}_{ik}}^{2}}\sqrt{\sum_{i}{\zeta^{Q}_{ik}}^{2}}}$
where $\zeta^{P}_{ik}=\eta^{P}_{ik}-\bar{\eta}^{P}_{k}$, $\eta^{P}_{ik}=-\frac{1}{\log 2}\left(P_{ik}\log P_{ik}+(1-P_{ik})\log(1-P_{ik})\right)$, $\bar{\eta}^{P}_{k}=\frac{1}{N}\sum_{i}\eta^{P}_{ik}$, and scalars $\zeta^{Q}_{ik}$, $\eta^{Q}_{ik}$, and $\bar{\eta}^{Q}_{k}$ are defined analogously.
While it is possible to define (hard) accuracy for this multilabel case, a more common evaluation metric is the micro F1 score, which in our notation can be expressed as:
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}2\frac{\sum_{ik}J^{P}_{ik}J^{Q}_{ik}}{\sum_{ik}\left(J^{P}_{ik}+J^{Q}_{ik}\right)}$
where $J^{P}_{ik}=\mathds{1}(P_{ik}>0.5)$, and $J^{Q}_{ik}$ is defined analogously. Similarly, we define the soft version of the micro F1 score as follows (see derivations in the Appendix):
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}2\frac{\sum_{ik}\min(P_{ik},Q_{ik})}{\sum_{ik}(P_{ik}+Q_{ik})}.$
The (hard) micro F1 score is a special case of this soft counterpart when $P_{ik}\in\{0,1\}$ and $Q_{ik}\in\{0,1\}$ (i.e. no label variation). Furthermore, the soft accuracy in Section˜3.1 is a special case of this soft micro F1 score when the unity constraints hold. Therefore, the soft micro F1 score is a general evaluation metric for classification tasks.
3.3 Relationship between Metrics
To analyse the relationship between metrics, we compute Pearson correlation between each pair of multiclass metrics (excluding the macro ones) similar to prior work (Chicco and Jurman, 2020). For a given number classes, we sample $B$ matrices $\mathbf{P}$ and $\mathbf{Q}$ and compute the value of each metric on these $B$ pairs of $\mathbf{P}$ and $\mathbf{Q}$. The Pearson correlations between pairs of metrics are then computed based on these $B$ data points. We set $B=500$ for computational reasons. We report the correlation coefficients in Table˜2.
Table˜2 shows that the metrics are generally weakly correlated with each other, with the exception of soft accuracy and PO-JSD where the two are consistently strongly correlated ($r>0.9$). This strong correlation with an established HLV metric gives an empirical assurance that our proposed soft accuracy is a sensible metric. The table also shows that soft accuracy is only moderately correlated with (hard) accuracy in most cases, where this correlation gets weaker when the number of classes is large ($K=100$), or either $\mathbf{P}$ or $\mathbf{Q}$ is dense ($\alpha=10$ or $\beta=10$). This finding suggests that hard and soft accuracy capture different aspects when humans disagree. Furthermore, the table shows that entropy correlation is poorly correlated with all metrics, which suggests that it is an outlier HLV metric.
Soft Accuracy and PO-JSD
Table 2: Pearson correlations between a pair of evaluation metrics $m(\mathbf{P},\mathbf{Q})$ for 1K examples and various number of classes ($K$) where the rows of $\mathbf{P}$ and $\mathbf{Q}$ are drawn from a symmetric Dirichlet with parameters $\alpha$ and $\beta$ respectively. A, J, E, and S denote the accuracy, PO-JSD, entropy correlation, and soft accuracy respectively.
$K$ A-J A-E A-S J-E J-S E-S
$\alpha=\beta=10$
10 0.272 3470.272\,347 0.050 2550.050\,255 0.289 9840.289\,984 0.035 9270.035\,927 0.942 2630.942\,263 0.100 3030.100\,303
100 0.047 5880.047\,588 0.040 9640.040\,964 0.051 9560.051\,956 0.091 1220.091\,122 0.924 1290.924\,129 0.087 5600.087\,560
$\alpha=\beta=0.1$
10 0.627 7730.627\,773 0.063 8200.063\,820 0.733 7430.733\,743 0.126 1640.126\,164 0.963 2140.963\,214 0.162 8160.162\,816
100 0.194 0690.194\,069 0.093 1130.093\,113 0.251 3550.251\,355 0.014 8690.014\,869 0.961 0860.961\,086 0.032 2250.032\,225
$\alpha=10,\beta=0.1$
10 0.226 1750.226\,175 0.015 882-0.015\,882 0.206 4640.206\,464 0.018 527-0.018\,527 0.975 1440.975\,144 0.002 286-0.002\,286
100 0.004 257-0.004\,257 0.062 204-0.062\,204 0.026 900-0.026\,900 0.050 0800.050\,080 0.957 8160.957\,816 0.061 6400.061\,640
$\alpha=0.1,\beta=10$
10 0.211 8740.211\,874 0.016 4670.016\,467 0.181 7060.181\,706 0.016 632-0.016\,632 0.975 8300.975\,830 0.019 553-0.019\,553
100 0.103 4780.103\,478 0.021 530-0.021\,530 0.057 2680.057\,268 0.005 990-0.005\,990 0.967 4200.967\,420 0.005 4470.005\,447
Despite their strong positive correlation, soft accuracy and PO-JSD have a notable difference: soft accuracy is upper-bounded by PO-JSD.
Theorem 3.1.
Consider the same definitions of $P_{ik}$ and $Q_{ik}$ used in Section˜3. The soft accuracy and the PO-JSD between $\mathbf{P}$ and $\mathbf{Q}$ satisfy the following inequality:
$\frac{1}{N}\sum_{ik}\min(P_{ik},Q_{ik})\leq 1-\frac{1}{N}\sum_{i}\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i})$
where $\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i})$ denote the Jensen-Shannon divergence between the $i$-th rows of $\mathbf{P}$ and $\mathbf{Q}$.
Proof 3.2.
Given a scalar $0\leq\pi\leq 1$ and discrete distributions $\mathbf{a},\mathbf{b}$, it has been shown that the following bound holds (Lin, 1991, Theorem 4):
$\sum_{k}\min(\pi a_{k},(1-\pi)b_{k})\leq\frac{1}{2}\left(H([\pi,1-\pi])-\operatorname{JSD}_{\pi}(\mathbf{a},\mathbf{b})\right)$
where $H(\mathbf{u})$ is the entropy of $\mathbf{u}$, and $\operatorname{JSD}_{\pi}(\mathbf{a},\mathbf{b})=H(\pi\mathbf{a}+(1-\pi)\mathbf{b})-\pi H(\mathbf{a})-(1-\pi)H(\mathbf{b})$ is the general form of Jensen-Shannon divergence with two distributions. It can be easily shown that $\operatorname{JSD}_{\pi}$ is reduced to the standard Jensen-Shannon divergence when $\pi=\frac{1}{2}$. Performing this substitution and noting that $H([\frac{1}{2},\frac{1}{2}])=1$, [^5] All logarithms are of base 2. we have
$\sum_{k}\min(\frac{1}{2}P_{ik},\frac{1}{2}Q_{ik})\leq\frac{1}{2}\left(1-\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i})\right)$
for all $i$. Doubling both sides and taking the average of this inequality over $i$ give the desired result.
This fact implies that soft accuracy penalises a model more heavily than PO-JSD, as illustrated in Figure˜1. It shows that a skewed predicted judgement distribution of $(0.2,0.8)$ has a soft accuracy of $0.7$, much less than the PO-JSD (>0.9>0.9). The high PO-JSD is potentially misleading as it is very close to the maximum possible value of 1.
4 Training Methods for HLV
Figure 1: Soft accuracy and PO-JSD graphs for a binary classification problem on a single example where the true and the predicted judgement for the positive class is 0.5 and $q$ respectively.
Table 3: Overview of HLV training methods used, along with details of whether annotator identity is required (I), multilabel tasks are supported (M), and training data is enlarged (D) which makes training run much longer.
Name I M D
Repeated labelling (ReL)
Majority voting (MV)
Annotator ranking (AR)
Annotator ranking (hard) (ARh)
Ambiguous labelling (AL)
Soft labelling (SL)
Multitask of SL and MV (SLMV)
Annotator ensemble (AE)
Annotator ensemble (hard) (AEh)
Below are proposed in this work
Jensen-Shannon divergence (JSD)
Soft micro F1 (SmF1)
Soft macro F1 (SMF1)
Loss aggregation with $\min$ (LA-min)
Loss aggregation with $\max$ (LA-max)
We now discuss the training methods. In total, we experiment with 14 HLV training methods, as summarised in Table˜3. The first method (ReL) is the simplest and uses standard cross-entropy loss as the training objective based on prior work (Sheng, Provost, and Ipeirotis, 2008). The next 8 methods (MV, AR, ARh, AL, SL, SLMV, AE, AEh) are from the Learning with Disagreements shared task (Leonardelli et al., 2023) which also use standard cross-entropy loss. The next 3 methods (JSD, SmF1, SMF1) use soft metrics that are differentiable as the training objective. The last 2 methods (LA-min, LA-max) aggregate the cross-entropy losses of all annotations for a given input. In detail:
  1. 1.
    ReL (repeated labelling) trains on the disaggregated labels directly (Sheng, Provost, and Ipeirotis, 2008). For example, if the training data only has a single instance $x$ with annotations $y_{1}$ and $y_{2}$ then ReL constructs new training data $\{(x,y_{1}),(x,y_{2})\}$.
  2. 2.
    MV trains on the majority-voted labels. This is the baseline model given by the shared task organisers.
  3. 3.
    AR (annotator ranking) and ARh (annotator ranking hard) weight each training instance equal to the sum of ranking scores of its annotators (Cui, 2023). Suppose that training instance $x_{1}$ is labelled as A, A, B, and A by 4 annotators; and $x_{2}$ is labelled as A, B, and B by only the first 3 annotators. Then, to compute the ranking score of an annotator:
    1. (a)
      AR computes the average judgement of the majority-voted label given by the annotator. The annotator ranking scores are (in order) $\frac{3}{4}$, $\frac{1}{2}\left(\frac{3}{4}+\frac{2}{3}\right)$, $\frac{2}{3}$, and $\frac{3}{4}$.
    2. (b)
      ARh computes the average number of times the annotator agrees with the majority. The scores are $\frac{1}{2}$, $\frac{2}{2}$, $\frac{1}{2}$, and $\frac{1}{1}$.
    For both methods, the ranking score is zero if the annotator never agrees with the majority. Given an instance, the training objective is maximising the probability of its majority-voted label(s), scaled by its weight as defined above.
  4. 4.
    AL (ambiguous labelling) labels ambiguous instances, i.e. instances where annotators are not unanimous in their judgement, with all selected classes exactly once (Gajewska, 2023). Each instance-label pair is included in the training set. For example, if training instance $x_{1}$ is labelled as $y_{1}$, $y_{1}$, and $y_{2}$ by 3 annotators, and $x_{2}$ is labelled as $y_{2}$, $y_{2}$, and $y_{2}$, then AL produces $\{(x_{1},y_{1}),(x_{1},y_{2}),(x_{2},y_{2})\}$ as the training data. Because it was developed for single-label tasks, we exclude this method from multilabel tasks. [^6] Extending this method to multilabel tasks is beyond the scope of this work.
  5. 5.
    SL (soft labelling) trains on soft labels as targets (Maity et al., 2023; Wan and Badillo-Urquiola, 2023). For example, if an instance is labelled as positive by 80% of annotators then the log-likelihood is $0.8\log p(1)+0.2\log p(0)$, where $p(1)$ and $p(0)$ denote the predicted probability of the positive and the negative class respectively.
  6. 6.
    SLMV performs multi-task learning using both the soft (SL) and the majority-voted (MV) labels as targets (Grötzinger, Heuschkel, and Drews, 2023). Two models with a shared encoder are trained to predict the two types of labels respectively, which is similar to the method of Fornaciari et al. (2021). For example, if an instance is labelled as positive by 80% of annotators then the multitask log-likelihood is $0.8\log p_{\theta_{1}}(1)+0.2\log p_{\theta_{1}}(0)+\log p_{\theta_{2}}(1)$, where $\theta_{1}$ and $\theta_{2}$ denote the parameters of the first and the second model respectively. We predict only the soft labels for evaluation.
  7. 7.
    AE (annotator ensembling) and AEh (annotator ensembling hard) model each annotator separately followed by ensembling (Sullivan, Yasin, and Jacobs, 2023; Vitsakis et al., 2023). In this approach, we have as many models as there are annotators. Each model is trained on the labels given by a distinct annotator. In practice, we share the encoder parameters across models. Suppose that there are 3 annotators, and the predicted judgements of the positive class for an instance are 60%, 30%, and 90%. To ensemble these predictions at test time:
    1. (a)
      AE computes a simple mean (Sullivan, Yasin, and Jacobs, 2023). The final predicted judgement for the positive class is $\frac{1}{3}\left(0.6+0.3+0.9\right)$.
    2. (b)
      AEh computes the distribution of most likely labels (Vitsakis et al., 2023). The final predicted judgement for the positive class is $\frac{2}{3}$ because 2 out of 3 judgements are over 50%.
  8. 8.
    JSD, SmF1 (soft micro F1), and SMF1 (soft macro F1) train using JSD, 11- soft accuracy for multiclass or soft micro F1 for multilabel tasks, and 11- soft macro F1 as the objective respectively (see Section˜3 for definitions). [^7] We exclude entropy correlation because it doesn’t have a unique maximiser. These approaches leverage the differentiability of soft metrics, resulting in a single objective for both training and inference. By directly optimising the evaluation metric at training time, we expect the models to exhibit superior performance, especially when evaluated with the corresponding metric.
  9. 9.
    LA-min (loss aggregation $\min$) and LA-max (loss aggregation $\max$) aggregate the losses of all annotations for a given training instance using the $\min$ and the $\max$ functions respectively. For example, given a training instance $x$ with annotations $y_{1}$ and $y_{2}$, we compute the losses $l_{1}=\mathcal{L}(x,y_{1})$ and $l_{2}=\mathcal{L}(x,y_{2})$ where $\mathcal{L}$ is the cross-entropy loss function. Then, we update model parameters based on the gradient of $g(l_{1},l_{2})$ where $g$ is either the $\min$ or the $\max$ function. Using the $\min$ (resp. $\max$) function means selecting the least (resp. most) “surprising” annotation for the model. Using the mean aggregation is mathematically equivalent to SL and thus excluded.
5 Evaluation of HLV Training Methods
We now conduct an extensive study covering 6 datasets, 14 training methods, and 6 evaluation metrics across 2 pretrained models to understand the performance landscape of HLV data. Table˜4 shows an overview of the 6 datasets. These datasets are selected to represent diverse factors such as language, type of annotators, number of annotations for each example, number of classes, presence of annotator identity, and type of classification tasks (i.e. single- vs multilabel). By using these diverse datasets, consistent patterns that emerge across datasets are more likely to hold true in general rather than just in a specific type of datasets.
5.1 Experimental Setup
Table 4: Datasets used in the evaluation of HLV training methods, along with the language (L), type of annotators (O) where E and C mean Experts and Crowds respectively, number of examples to the nearest thousand (NN), average number of annotations per example ($J$), number of classes ($K$), whether annotator identity is present (I), and whether the task is multilabel classification (M).
Name Task L O NN $J$ $K$ I M
HS-Brexit (Akhtar, Basile, and Patti, 2021) Hate speech EN E 1 $6.0$ 2
MD-Agreement (Leonardelli et al., 2021) Offensiveness EN C 10 $5.0$ 2
ArMIS (Almanea and Poesio, 2022) Misogyny AR E 1 $3.0$ 2
ChaosNLI (Nie, Zhou, and Bansal, 2020) NLI EN C 3 $100.0$ 3
MFRC (Trager et al., 2022) Moral sentiment EN E 18 $3.4$ 8
TAG (private dataset) Legal area EN E 11 $5.5$ 33
Public datasets
We use 5 public datasets with disaggregated annotations that have been used in previous HLV studies, four of which are in English and one in Arabic. For HS-Brexit, MD-Agreement, and ArMIS, we use the same train–test splits as Leonardelli et al. (2023). For both ChaosNLI and MFRC, we use 10-fold cross-validation with 10% of the training portion used as a development set. For ChaosNLI, we include only the SNLI and MNLI subsets (each has 1.5K examples) because of their standard NLI setup of one premise and one hypothesis.
“Text Annotation Game” (TAG) dataset
We also include a private dataset called TAG which consists of English texts describing legal problems written by non-expert legal help seekers. Because it is based on real-world confidential legal requests from help-seekers, we are unable to distribute the dataset. However, we still include the TAG dataset due to its importance for the meta-evaluation later in Section˜6. We have access to this dataset because of our collaboration with Justice Connect, [^8] https://justiceconnect.org.au an Australian public benevolent institution [^9] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile providing free legal assistance to laypeople facing legal problems. Each text in the dataset was annotated by an average of 5.5 practising lawyers, who each selected one or more out of the 33 areas of law [^10] Including a special label “Not a legal issue”. that applied to the problem. Thus, it is a multilabel classification task. Example areas of law include Neighbourhood disputes and Housing and residential tenancies. Average inter-annotator agreement ($\alpha$) over the areas of law is $0.454$, which is modest. [^11] We compute this agreement on random 10% selection of the data due to the high computational cost. This is because lawyers often have different interpretations of the same problem due to their different legal specialisations and years of experience. However, these different interpretations are all valid as human label variation because these registered lawyers are subject matter experts who play a crucial part in interpreting the law. The dataset has a total of 11K texts collected between July 2020 and early December 2023. We randomly split the dataset 8:1:1 for the training, development, and test sets respectively.
Models
We experiment with both small and large language models. Because HS-Brexit, MD-Agreement, and ArMIS use Twitter as source data, we use TwHIN-BERT (Zhang et al., 2023) which was pretrained on Twitter data and supports both English and Arabic. For other datasets, we use RoBERTa (Liu et al., 2019). In addition, we experiment with 8B LLaMA 3 (Grattafiori et al., 2024) for the English datasets. [^12] While LLaMA 3 was trained on multilingual data, its intended use is English only. We replace the output layer of these pretrained models with a linear layer with $K$ output units. We use the base versions of both RoBERTa and TwHIN-BERT (with roughly 100M parameters each).
Training
We implement all methods using FlairNLP (Akbik et al., 2019). For RoBERTa and TwHIN-BERT models, we tune the learning rate and the batch size using random search. Because there are multiple evaluation metrics that aren’t necessarily comparable, we select the best hyperparameters based on the geometric mean of their values [^13] We transform entropy correlation to a value between 0 and 1 before taking the mean. to avoid biasing towards one metric. We exclude both hard and soft macro F1 scores from this mean as most datasets are balanced. For LLaMA, we use FlairNLP’s default hyperparameters [^14] Learning rate and batch size are $5\text{\times}{10}^{-5}$ and 32 respectively. for computational reasons. We use LoRA (Hu et al., 2022) to finetune LLaMA efficiently. All models are finetuned for 10 epochs. We truncate inputs longer than 512 tokens in the TAG dataset, affecting only 4% of instances. For ChaosNLI and MFRC, we train the final model used for evaluation of each fold on the concatenation of the training and the development sets.
Evaluation
When cross-validation is used, we evaluate the concatenated test predictions. Otherwise, we evaluate on the test set. We perform the evaluation across 3 training runs.
5.2 Results
Figure 2: Performance difference (delta) with mean MV performance of TwHIN-BERT on ArMIS. The methods are sorted by their mean ranking across datasets, models, and metrics.
Figure 3: Performance difference (delta) with mean MV performance on HS-Brexit and MD-Agreement. SmF1 with TwHIN-BERT predicts zero for all test instances on HS-Brexit so its entropy correlation is undefined. The methods are sorted by their mean ranking across datasets, models, and metrics.
Figure 4: Performance difference (delta) with mean MV performance on ChaosNLI (both the SNLI and the MNLI portions), MFRC, and TAG datasets. ChaosNLI doesn’t have annotator identity information, so AE, AEh, AR, ARh are inapplicable. AL is omitted from MFRC and TAG because it is developed for single-label rather than multilabel tasks. AEh with RoBERTa predicts zero for all test instances on several classes in TAG, so its entropy correlation is undefined. The methods are sorted by their mean ranking across datasets, models, and metrics.
We compute the difference between the performance of a single run of an HLV training method and the mean performance of MV. We then report the mean differences across runs. We present the results on ArMIS in Figure˜2 (the only non-English dataset), HS-Brexit and MD-Agreement in Figure˜3 (Learning with Disagreements shared task datasets), and the remaining datasets in Figure˜4 (multiclass and multilabel datasets). We report performance of MV in Table˜7 in the Appendix. We make the following observations.
First, HLV training methods have different effectiveness compared to MV, with some methods performing very poorly. For example, on ArMIS and both portions of ChaosNLI, LA-min consistently performs worse than MV. This finding suggests the importance of choosing the right training method when embracing HLV.
Second, both ReL and SL substantially outperform MV in 68 out of 78 settings (87.2% of all settings), [^15] Negative cases for ReL: all metrics except PO-JSD on ArMIS (5 settings), LLaMA on HS-Brexit and MD-Agreement with accuracy (2), LLaMA on MD-Agreement with macro F1 score (1), LLaMA on TAG with hard metrics (2). Negative cases for SL: TwHIN-BERT on ArMIS with hard metrics (2), LLaMA on HS-Brexit with all metrics except entropy correlation and soft macro F1 score (4), LLaMA on MD-Agreement and TAG with hard metrics (4). where a setting is a 3-tuple of a dataset, a model, and an evaluation metric. [^16] We consider the SNLI and the MNLI portions of ChaosNLI as separate datasets for this purpose. Furthermore, ReL or SL is among the top-performing methods in 89.7% of all settings, [^17] Negative cases: TwHIN-BERT on HS-Brexit with accuracy (1 setting); RoBERTa on MFRC with entropy correlation, PO-JSD, and soft macro F1 score (3); LLaMA on MFRC with macro F1 score, PO-JSD, and both micro and macro soft F1 scores (4). including when compared to JSD and SmF1 evaluated using PO-JSD and soft accuracy/micro F1 respectively. Concretely, ReL and SL are on par with (or outperform in some cases with LLaMA) the training method whose learning objective is the same as the evaluation metric in 23 out of 26 settings (88.5%), [^18] We consider only PO-JSD and soft accuracy/micro F1 evaluation metrics to get the total of 26 settings. Negative cases: RoBERTa on MFRC (2 settings), LLaMA on MFRC with PO-JSD (1). demonstrating the strength of ReL and SL.
Third, ReL and SL are generally competitive with each other. However, there are two exceptions to this trend. First, on ArMIS, SL outperforms ReL. Second, on both HS-Brexit and ChaosNLI, ReL outperforms SL with LLaMA. For the former finding, however, ArMIS is an outlier dataset: it is the only Arabic dataset, has the smallest number of examples and annotations per example, and all methods generally have higher variance on this dataset. The latter finding suggests that ReL is better suited to large language models than SL.
Fourth, SMF1 and JSD generally also perform well, although they lag behind ReL and SL. Because both SMF1 and JSD use a soft evaluation metric as the training objective, this finding underlines the value in using differentiable metrics for HLV model training.
We also observe that SmF1 with TwHIN-BERT results in undefined entropy correlation on HS-Brexit. Looking closely, we find that the method degenerates to predicting $Q_{ik}=0$ for all $i$. We observe the same degenerate results for AEh with RoBERTa on the TAG dataset. Moreover, SmF1 with RoBERTa attains much lower soft micro F1 score than the top-performing ReL on TAG, even though SmF1 optimises the evaluation metric directly during training. Our investigation suggests that these unexpected observations are due to the class imbalance present in both HS-Brexit and TAG. When the training objective is fair (e.g., SMF1), the problem disappears.
5.3 Discussion
ReL and SL appear to perform best as a training method, and that is somewhat surprising because they are the most straightforward methods to incorporate HLV in model training. That said, our finding that SL is a strong performer is in line with what Uma et al. (2021) discovered. The results of ReL, however, stand in contrast with prior work (Uma et al., 2021; Kurniawan et al., 2024). This difference may be due to the structured prediction tasks that the prior work considered which have an exponentially large output space. We note that Uma et al. (2021) also found that ReL performs really well in simple, multiclass classification tasks, in line with our findings.
Moreover, the results suggest that ReL is likely to outperform SL when large language models are used. A possible explanation is that with SL, the model observes the complete judgement distribution of an instance as target. Thus, there is not much flexibility as the model has to predict that distribution to minimise the training loss. In contrast, with ReL, the model only observes one annotation of the instance as target. Therefore, it has more flexibility in how it distributes the judgement distribution mass across all the labels of that instance over training steps. We hypothesise that large language models have sufficient learning capacity to benefit from this flexibility while smaller models do not. This explanation is consistent with our findings on ChaosNLI that ReL has much larger gains compared to SL. ChaosNLI has 100 annotations per instance, the largest number out of all datasets. We leave further investigation of this hypothesis to future work.
Despite its strengths, ReL has substantial computational cost: because it keeps the annotations disaggregated, the size of the training data can be enormous. For example, in the ChaosNLI dataset, ReL makes the training data 100 times larger because there are exactly 100 annotations for each training instance. That ReL performs better with LLaMA than RoBERTa on ChaosNLI exacerbates the computational inefficiency of ReL: both the data and the model must be large for it to perform optimally. Mitigating this inefficiency can be an interesting avenue for future work.
Because the training data is much larger, one may think that the superiority of ReL over SL is due to its much longer training. However, this is not the case. We find that increasing the number of training iterations of SL to match that of ReL does not improve performance.
6 Meta-Evaluation of HLV Evaluation Metrics
In the previous section, we saw the best training methods that perform consistently well across metrics. In this section, we address the next question: what is the best evaluation metric for HLV data? We answer this by conducting a meta-evaluation on the evaluation metrics. We focus on the TAG dataset specifically for this meta-evaluation experiment as we have access to practicing lawyers as annotators who can provide high quality annotations. It does mean we are unable to distribute the dataset for ethical reasons, but we believe the insights derived from the meta-evaluation is worth the trade-off.
A good evaluation metric should produce a ranking of training methods (henceforth “metric ranking”) that correlates to the ranking produced by human judgements (henceforth “human ranking”). To create the latter (human ranking), we frame it as a pairwise comparison task where we pit two methods against each other and ask lawyers to select one of them. Given a large number of human-judged pairwise comparisons, we then use an algorithm to create a ranking for the training methods. Intuitively, the algorithm will rank a method that “wins” consistently higher than another method that “loses” most of the time. Once we have the human ranking, we can then compute its correlation with each metric ranking, and the best metrics are the ones with the highest correlations.
6.1 Experimental Setup
[…]
To make sure the lawyers form their own expectation on the areas of law relevant to an instance before making this pairwise selection, we first show the instance text and ask them to select the most relevant areas of law (Figure˜6 in the Appendix). After completing this task for an instance text, we then ask them to make the pairwise selection. The two tasks must be completed for an instance text before the lawyers can move to the next instance text. We conduct several pilot studies with the lawyers to refine the interface design and ensure the task is clear to them.
At the end of the annotation task, we have a collection of human-judged pairwise comparisons for the training methods. We next use the rank centrality algorithm (Negahban, Oh, and Shah, 2012) to compute the “score” for each method to ultimately produce a ranking of the methods. This algorithm casts the problem as a random walk on a weighted directed graph where each node corresponds to a method and the weight of the edge from method $i$ to method $j$ is the probability that method $j$ wins over model ii. [^21] We determine that method $j$ wins over method $i$ if the annotator select method $j$ but not method $i$. The score of a method is then its stationary probability for this random walk. Intuitively, the score is high if it wins against other high-scoring methods or against many low-scoring methods. To assess the evaluation metrics and understand which metrics are the best, for each metric we compute the Pearson correlation between: (1) method performance as given by the evaluation metric; and (2) the method scores produced by human judgements.
6.2 Results
We draw the regression plots between method performance as given by the evaluation metrics and method scores produced by human judgements in Figure˜5. The figure shows that not all metrics exhibit a strong linear relationship with human judgements. For example, both hard F1 scores and entropy correlation show almost no relationship for LLaMA. All metrics except entropy correlation show a strong linear relationship for RoBERTa. This suggests that entropy correlation is a poor HLV metric.
In Table˜5, we show the Pearson correlation coefficients for all metrics along with the $p$-values. The table confirms that all metrics are positively correlated with human judgements but with different magnitudes. As before, both hard F1 scores and entropy correlation are very weakly correlated with human judgements for LLaMA, and entropy correlation is the weakest for RoBERTa. Furthermore, it is the only metric whose correlation is not statistically significant. Across both pretrained models, both PO-JSD and soft micro F1 consistently have the strongest correlation, but each is best for different models: PO-JSD is best for LLaMA while soft micro F1 score is best for RoBERTa. This finding suggests that both metrics are good for evaluation in the HLV context.
6.3 Method ranking based on human judgements
Given the scores computed from human judgements, we can rank the methods based on their scores (human ranking). Table˜6 shows that the human rankings have some similarities to TAG results in Figure˜4. For instance, ReL, JSD and SL are generally strong performers (exception: ReL with LLaMA). That said, we also see some discrepancies. For example, SLMV is highly ranked here but not reflected in Figure˜4. We contend, however, that the meta-evaluation study is ultimately based on one dataset and so we should interpret these results cautiously.
Figure 5: Relationships between method performance as given by the evaluation metrics (Value) and method scores produced by human judgements (Score) on the TAG dataset. The shaded area denotes a 95% confidence interval. Each data point corresponds to a specific training method (e.g., ReL). AEh is excluded from RoBERTa with entropy correlation because of its degenerate results.
Table 5: Correlation coefficients between method performance as given by the evaluation metrics and method scores produced by human judgements on the TAG dataset along with their $p$-values computed with a permutation test. Asterisks denote statistical significance ($p<.05$).
Model Type Metric Pearson $r$ $p$-value
RoBERTa Hard Micro F1 $0.787$ $0.001$
Macro F1 $0.774$ $0.005$
Soft Entropy correlation $0.339$ $0.281$
PO-JSD $0.674$ $0.005$
Soft micro F1 $0.804$ $0.001$
Soft macro F1 $0.753$ $0.006$
LLaMA Hard Micro F1 $0.002$ $0.984$
Macro F1 $0.087$ $0.776$
Soft Entropy correlation $0.125$ $0.689$
PO-JSD $0.539$ $0.070$
Soft micro F1 $0.343$ $0.249$
Soft macro F1 $0.213$ $0.488$
Table 6: Rankings of HLV training methods based on their scores produced by human judgements on the TAG dataset.
RoBERTa LLaMA
Rank 1–7 Rank 8–13 Rank 1–7 Rank 8–13
Method Score Method Score Method Score Method Score
SMF1 $0.126$ MV $0.076$ SL $0.130$ ReL $0.070$
JSD $0.109$ SmF1 $0.070$ SLMV $0.117$ LA-max $0.061$
ARh $0.102$ LA-max $0.061$ AE $0.105$ ARh $0.055$
SLMV $0.098$ AR $0.048$ MV $0.092$ LA-min $0.054$
SL $0.095$ AE $0.027$ AR $0.088$ AEh $0.050$
ReL $0.085$ AEh $0.024$ JSD $0.086$ SMF1 $0.014$
LA-min $0.078$ SmF1 $0.076$
6.4 Discussion
Putting all the results together, PO-JSD and soft micro F1 are arguably some of the best metrics for HLV data. That said, Uma et al. (2021) argued that the value of PO-JSD is typically confined within a small range and lacks an intuitive interpretation. We believe that this is its biggest weakness, which is shared by other information-theoretic measures such as cross-entropy. In contrast, soft micro F1 is more interpretable as it is analogous to the standard F1 score. Therefore, our general recommendation is to report both metrics, but focus on the soft micro F1 score when an accessible interpretation is important (e.g., communicating with non-technical people) or one is restricted to a single metric (e.g., in hyperparameter tuning).
7 Conclusions
We propose new evaluation metrics for evaluating model predictions with human label variation (HLV). Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. Therefore, our proposed metrics have intuitive interpretations and reduce to standard hard metrics if there is no label variation. While our analysis shows that our proposed soft accuracy metric is strongly correlated with an existing metric based on Jensen-Shannon divergence, we mathematically prove that the former is upper-bounded by the latter. We further show that the JSD-based metric can produce a misleadingly high score as a result.
Because the soft metrics are differentiable, we propose 3 new training methods for HLV using the metrics as the training objective. Additionally, we propose 2 new training methods that aggregate losses over annotations of the same instance, bringing the total of our proposed training methods to 5 methods. We test our proposed methods on 6 datasets spanning binary, multiclass, and multilabel classification tasks, as well as crowd and expert annotators. We evaluate against 9 existing HLV training methods across 2 pretrained models and use a total of 6 HLV evaluation metrics including both existing and proposed ones to find the best methods for HLV data. Then, we perform an empirical meta-evaluation of the evaluation metrics to understand which metrics are best for HLV data.
We find that simple methods such as training on disaggregated annotations (ReL) or soft labels (SL) perform best in most cases. They often outperform not only training on the majority-voted labels but also more complex HLV training methods including our proposed training methods with the differentiable metrics. Our meta-evaluation shows that our proposed soft micro F1 score is one of the best metrics for HLV data. This metric reduces to our proposed soft accuracy in single-label classification. Given its intuitive interpretation and positive meta-evaluation result, we recommend further research to include soft micro F1 when reporting model performance in the HLV context.
Limitations
The TAG dataset is private and cannot be released publicly, which is a limitation of our work in terms of reproducibility. This is because the data is based on real-world confidential legal requests from help-seekers. More importantly, their safety and privacy is of high concern: some cases are so unique that even if you were to anonymise the cases, re-identfication may still be possible. Nevertheless, we believe our work still offers valuable scientific knowledge on the investigation of HLV training methods and evaluation metrics.
Our empirical meta-evaluation is limited to a single multilabel classification dataset. Thus, it is unclear if the findings extend to other datasets and binary or single-label tasks. Future work could expand on this limitation by experimenting with a different or diverse selection of datasets.
[…]
Derivation of Soft F1 Scores
.1 Fuzzy Sets
Below we provide a brief overview of fuzzy sets drawn from the work of Kruse et al. (2022). A fuzzy set $A$ is defined by its membership function $\mu_{A}:X\to[0,1]$, where $\mu_{A}(x)$ represents the degree of membership of $x\in X$ in $A$. The cardinality of $A$ is defined as the sum of the degrees of membership of all elements of $X$:
$\sum_{x\in X}\mu_{A}(x).$
If $A$ and $B$ are fuzzy sets over the same universe $X$ then their intersection is a fuzzy set whose membership function is defined as
$\mu_{A\cap B}(x)=t(\mu_{A}(x),\mu_{B}(x))$
[…]
.2 Soft F1 Scores
Let $H_{k}$ and $R_{k}$ denote the predicted and the true (crisp) sets of examples in class $k$ respectively. Standard precision and recall scores for class $k$ are defined as:
$\mathrm{Prec}_{k}$ $=\frac{|H_{k}\cap R_{k}|}{|H_{k}|},$
$\mathrm{Rec}_{k}$ $=\frac{|H_{k}\cap R_{k}|}{|R_{k}|}.$
Embracing HLV, we view judgement distributions as degrees of memberships of fuzzy sets, each corresponds to a class, over a universe of examples. Specifically, we consider $P_{ik}$ and $Q_{ik}$ as the true and the predicted degrees of membership of example $i$ in class $k$. Therefore, we can define the soft precision and recall scores using fuzzy set operations as follows (Harju and Mesaros, 2023):
$\mathrm{SoftPrec}_{k}=\frac{\sum_{i}\min(P_{ik},Q_{ik})}{\sum_{i}Q_{ik}}$ (1)
and
$\mathrm{SoftRec}_{k}=\frac{\sum_{i}\min(P_{ik},Q_{ik})}{\sum_{i}P_{ik}}.$ (2)
The soft F1 score is the harmonic mean between the soft precision and recall scores as usual:
$\mathrm{SoftF\textsubscript{1}}_{k}=2\frac{\sum_{i}\min(P_{ik},Q_{ik})}{\sum_{i}(P_{ik}+Q_{ik})}.$ (3)
Taking the mean of Equation˜3 over classes results in the soft macro F1 score. To obtain the micro variant, we simply modify the sums in both the numerator and the denominator of Equations˜1 and 2 to also iterate over classes to get the soft micro precision and recall scores, and then take the harmonic mean of the two scores as normal.
[…]
Performance of MV
Table˜7 reports the performance of MV across datasets and evaluation metrics.
Table 7: Mean ($\pm$ std) performance of MV measured by entropy correlation, PO-JSD, and various versions of F1 scores. MD-Agr refers to MD-Agreement. C-SNLI and C-MNLI refer to the SNLI and the MNLI portions of ChaosNLI respectively. Hard (resp. soft) accuracy scores (for multiclass tasks) are reported in the hard (resp. soft) micro F1 score column.
Dataset Model Micro F1 Macro F1 Entropy corr. PO-JSD Soft micro F1 Soft macro F1
HS-Brexit TwHIN-BERT .903 $\pm$ .009 .581 $\pm$ .093 .470 $\pm$ .119 .941 $\pm$ .010 .883 $\pm$ .013 .676 $\pm$ .071
LLaMA .950 $\pm$ .012 .660 $\pm$ .104 .330 $\pm$ .065 .939 $\pm$ .002 .886 $\pm$ .003 .623 $\pm$ .012
MD-Agr TwHIN-BERT .809 $\pm$ .003 .779 $\pm$ .002 .379 $\pm$ .002 .921 $\pm$ .001 .811 $\pm$ .002 .790 $\pm$ .003
LLaMA .810 $\pm$ .001 .784 $\pm$ .001 .215 $\pm$ .019 .863 $\pm$ .003 .770 $\pm$ .002 .743 $\pm$ .003
ArMIS TwHIN-BERT .713 $\pm$ .014 .705 $\pm$ .014 .062 $\pm$ .015 .769 $\pm$ .011 .702 $\pm$ .009 .695 $\pm$ .008
C-SNLI RoBERTa .632 $\pm$ .005 .551 $\pm$ .007 .123 $\pm$ .014 .830 $\pm$ .005 .672 $\pm$ .004 .629 $\pm$ .007
LLaMA .601 $\pm$ .010 .551 $\pm$ .008 .118 $\pm$ .022 .830 $\pm$ .005 .668 $\pm$ .005 .633 $\pm$ .007
C-MNLI RoBERTa .496 $\pm$ .014 .369 $\pm$ .014 .044 $\pm$ .007 .854 $\pm$ .009 .668 $\pm$ .010 .642 $\pm$ .009
LLaMA .514 $\pm$ .017 .439 $\pm$ .010 .100 $\pm$ .018 .827 $\pm$ .005 .643 $\pm$ .006 .615 $\pm$ .006
MFRC RoBERTa .657 $\pm$ .001 .456 $\pm$ .002 .404 $\pm$ .002 .954 $\pm$ .000 .591 $\pm$ .001 .418 $\pm$ .000
LLaMA .649 $\pm$ .001 .454 $\pm$ .005 .353 $\pm$ .005 .948 $\pm$ .000 .576 $\pm$ .001 .395 $\pm$ .001
TAG RoBERTa .742 $\pm$ .002 .610 $\pm$ .014 .394 $\pm$ .002 .985 $\pm$ .000 .645 $\pm$ .002 .471 $\pm$ .003
LLaMA .759 $\pm$ .006 .681 $\pm$ .005 .279 $\pm$ .010 .983 $\pm$ .000 .648 $\pm$ .002 .521 $\pm$ .003
[…]
Annotation Interface Examples
Figures˜6 and 7 show the interface of the area of law and the preference annotation tasks respectively.
Figure 6: Areas of law annotation interface
Figure 7: Preference annotation interface
Acknowledgements.
This research is supported by the Australian Research Council Linkage Project LP210200917232323https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution. [^24] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.